### v1

# HSSBench: Benchmarking Humanities and Social Sciences Ability for Multimodal Large Language Models

Abstract
Multimodal Large Language Models (MLLMs) have demonstrated significant potential to advance a broad range of domains. However, current benchmarks for evaluating MLLMs primarily emphasize general knowledge and vertical step-by-step reasoning typical of STEM disciplines, while overlooking the distinct needs and potential of the Humanities and Social Sciences (HSS). Tasks in the HSS domain require more horizontal, interdisciplinary thinking and a deep integration of knowledge across related fields, which presents unique challenges for MLLMs, particularly in linking abstract concepts with corresponding visual representations. Addressing this gap, we present HSSBench, a dedicated benchmark designed to assess the capabilities of MLLMs on HSS tasks in multiple languages, including the six official languages of the United Nations. We also introduce a novel data generation pipeline tailored for HSS scenarios, in which multiple domain experts and automated agents collaborate to generate and iteratively refine each sample. HSSBench contains over 13,000 meticulously designed samples, covering six key categories. We benchmark more than 20 mainstream MLLMs on HSSBench and demonstrate that it poses significant challenges even for state-of-the-art models. We hope that this benchmark will inspire further research into enhancing the cross-disciplinary reasoning abilities of MLLMs, especially their capacity to internalize and connect knowledge across fields.
Figure 1: We propose HSSBench, a large-scale benchmark spanning six diverse categories and 45 types, comprising 13,152 samples collected in the six official languages of the United Nations.
11footnotetext: HSSBench is publicly available at: https://github.com/Zhaolu-K/HSSBench.
1 Introduction
Multimodal large language models (MLLMs) achiam2023gpt ; team2024gemini have demonstrated remarkable performance across a wide range of tasks, recently achieving or even exceeding human-level capabilities in many of them. As the performance of MLLMs continues to improve, conducting a comprehensive evaluation of their capabilities has become increasingly essential. In recent times, numerous benchmarks for evaluating MLLMs have emerged hendrycks2020measuring ; yue2024mmmu ; liu2024mmbench ; saikh2022scienceqa ; zhang2024cmmmu . These tasks are designed to assess the models’ ability to jointly understand and reason across multiple modalities, such as images and text, from various perspectives.
Specifically, most multimodal benchmarks are designed either from a general perspective liu2024mmbench ; liu2023visualinstructiontuning ; meng2024vgavisionguiassistant ; han2024instinctivebiasspuriousimages ; qian2024easyfoolmultimodalllms or with a focus on scientific disciplines such as mathematics wang2024measuringmultimodalmathematicalreasoning ; lu2024mathvistaevaluatingmathematicalreasoning science li2023scigraphqalargescalesyntheticmultiturn ; liang2024scemqascientificcollegeentrance , and programming song2025csbench . Unlike STEM fields that employ "vertical reasoning": a focused sequential process using logical deduction and experimental analysis to arrive at singular correct answers, the Humanities and Social Sciences (HSS) emphasize "horizontal reasoning," requiring connections across different contexts and generating multiple valid interpretations rather than single solutions. This fundamental difference stems from the inherent attributes of these disciplines: While STEM fields utilize relatively fixed symbolic systems with standardized reading sequences, HSS disciplines feature symbol systems deeply rooted in regional cultures with meanings that require historical-cultural context interpretation. Furthermore, STEM knowledge can be iteratively developed through logical deduction and experimental analysis, whereas HSS knowledge verification relies on more complex pathways involving cross-referencing literature and expert consensus, with strong dependencies on real-world information.
Although some efforts have been made to explore aspects of the HSS, these attempts lack depth and do not provide a comprehensive and thoughtful examination of MLLMs within the context of these fields. In the study of HSS-related problems, achieving cross-modal knowledge transfer is crucial. Take the scenario illustrated in Figure 2 as an example: humans with basic knowledge can accurately infer "Business Penmanship" from the image content. When directly asked about knowledge points related to "Business Penmanship", the MLLM provides correct answers. However, when queried indirectly through an image, the model fails to recognize the font features in the text, preventing it from associating the relevant knowledge points with the image. This reveals a problem: most models struggle to establish meaningful mapping relationships between HSS-related images and the abstract concepts they represent. Although these models may recognize abstract concepts in isolation, they fail to effectively connect these concepts with the corresponding visual content. From a long-term perspective, a model skilled at solving mathematical problems but unable to interpret historical contexts or understand ethical principles offers an incomplete and potentially harmful form of intelligence.
To address this challenge, we introduce HSSBench, an innovative and comprehensive multilingual benchmark specifically crafted to thoroughly assess the performance of MLLMs in the HSS domain. Comprising around 13k carefully curated test items, HSSBench is structured into 45 types and covers 6 key categories within the field. HSSBench utilizes the visual question answering (VQA) format for this purpose, as shown in Figure 1. Given the involvement of numerous key domains, our work requires interdisciplinary collaboration. To this end, we have engaged experts from various fields to design the data framework and ensure quality control, thereby maximizing the representativeness and credibility of HSS-related issues. In addition to domain experts contributing data, we leveraged the expertise of both specialists and MLLMs to develop a VQA generation pipeline, which produced a portion of the high-quality data. Finally, through evaluations supporting English, Chinese, French, Russian, Spanish, and Arabic, HSSBench enables the assessment of the capabilities of MLLMs in addressing HSS challenges across a wide range of linguistic contexts.
In this study, we evaluate the performance of HSSBench across a range of MLLMs and find that it presents a significant challenge for these models, as their accuracy often falls below 60%. We conducted several comparative experiments to analyze the performance of the models. Our contributions can be summarized as follows:
Figure 2: An example of cross-modal knowledge transfer issues in MLLMs within the HSS domain. They struggle to associate "Business Penmanship" font knowledge with relevant images or recognize fonts in image text.
  • First, we introduced HSSBench, a novel dataset specifically designed for the HSS domain, which encompasses 6 distinct categories and 45 major types of HSS tasks.
  • Second, we offered a practical data construction method. It utilizes a multi-agent framework tailored for the HSS domain, allowing batch generation of high-quality, novel datasets.
  • Finally, we conducted detailed evaluations of over 20 MLLMs on HSSBench across six languages, verifying that HSS tasks still pose significant challenges for MLLMs. This work establishes a foundation for future MLLM research focusing on HSS and serves as a benchmark for future studies in this field.
2 HSSBench
The worldview within the HSS is expansive and lacks unified definitions. Our dataset focuses on six high-interest categories within the HSS domain: geography, art, culture, social sciences, history, and economics. Textual data are the primary medium for disseminating knowledge in the HSS domain, with almost no data in pictorial form. This presents significant challenges for experts in the construction of datasets. For this purpose, we designed and developed a VQA Generation Pipeline (VGP) to generate the dataset. After engaging experts for annotation, we designed a data construction agent based on their annotation logic, enabling us to generate a sufficient volume of data.
In this section, we will outline the details of the VGP. We have divided this process into three main stages: Dataset Preparation, Dataset Construction, and Validation. The three stages are summarized in Figure 3. 1) During the Dataset Preparation stage, both experts and a networked agent participate in obtaining the raw materials necessary for data construction. Either source is sufficient to provide the foundational content for the subsequent stage. 2) In the Dataset Construction stage, data can be constructed using two methods: expert-constructed and agents-constructed approach. 3) Within the Validation stage, both experts and agent filtration work jointly to identify high-quality data. The screened data passes through the second stage again until it meets the required standards. Data that do not conform to the construction logic are removed.
2.1 Dataset Preparation, Stage \@slowromancapi@
At this stage, we focus on collecting data on target disciplines, covering texts and illustrations. Through systematic organization and strict screening, we lay a solid data foundation for the subsequent development of high-quality questions and answers.
Figure 3: Our pipeline to build HSSBench.
Data Selection and Collection with Experts
During the data preparation for HSSBench, ensuring data diversity is crucial. For image and text selection, we adopt a combined approach involving experts and a multi-agent framework.
For images, to avoid data leakage, we initially encourage experts from various fields to use their private images. When acquisition is difficult, the dataset also incorporates images drawn by experts and licensed ones from open-source communities obtained by both experts and the multi-agent framework. Experts and agents collaborate to screen image content, choosing those related to subject areas and targeted knowledge points for constructing high-quality questions.
Regarding text, as experts may face challenges like unfamiliarity with domain-specific knowledge, we recommend they acquire high-quality textual materials from multiple sources, such as academic resource repositories of universities and open-source communities. This enriches information sources and mitigates bias. In the early stages, domain experts obtain credible resources like textbooks, past papers, and digital course materials from various disciplines. Since these resources are proofread during compilation, they possess high reliability and dense knowledge content, making them trustworthy sources for constructing questions. Each expert reviews materials in their field, extracting relevant text passages and images, eliminating redundant information and standardizing the data format.
Networked Information-Aggregation Agent
Inspired by the expert-led data collection process, we designed the Networked Information-Aggregation Agent to mimic their workflow. This agent broadens data sources by filtering Internet data while maintaining quality. For each discipline, it first compiles relevant keywords as knowledge point indices. Then, it retrieves online data, classifies it into text and images, and matches data against the keywords to assess relevance. For text, it evaluates aspects like professionalism, uniqueness, logical structure, and the need for cross-validation with images. If the text meets criteria, related images are extracted. Through this process, we obtain high-quality multimodal data at a controlled cost. Finally, domain experts review the curated data to ensure professionalism and reliability.
2.2 Dataset Construction, Stage \@slowromancapii@
At this stage, we developed a multidisciplinary visual-question generation pipeline comprising two key phases: Experts Construction and Multi-Agent Construction. These phases are designed to ensure that the questions produced offer a comprehensive assessment of the model’s effectiveness.
Dataset Construction with experts.
We require each question to be presented in a multiple-choice format, with one correct answer and three distractors that were plausible but ultimately incorrect. At this stage, experts have two main responsibilities: revising existing questions and creating entirely new ones. 1) When revising existing questions, experts optimize multiple-choice items, encompassing both the original ones provided in the materials and those requiring revision after failing validation tests. Experts carefully review the relevant textual and visual content, adjusting the question stems and answer choices by integrating real-world knowledge. The goal is to strengthen the connection between the questions and the accompanying images while improving the plausibility of the incorrect options. 2) For creating new questions, experts generate items based on the given high-quality texts and images. These new questions are expected to demand identification of knowledge points from the images and complex reasoning based on real-world understanding to arrive at the correct answers. Following these principles, experts produce a set of high-quality questions that fully utilize the provided multimodal materials to meet the demands of model performance evaluation.
Dataset Construction with Multi-Agent.
Inspired by the expert annotation process and the high-quality questions it produces, we designed a multi-agent automated construction framework aimed at improving data generation efficiency and reducing human labor. This agent architecture comprises several roles, including summarizer, extractor, question generator, and image matcher.
Initially, the summarizer and extractor independently analyze the document text to produce a comprehensive summary and a set of high-quality text segments, respectively. The summary provides an overview of the key knowledge points within the document, while the extracted segments offer detailed information suitable for direct question formulation. LLM then scores these text segments based on information density, uniqueness, and logical coherence, selecting the top N segments according to their scores.
The question generator then formulates N questions by leveraging both the full summary and the selected text segments. It operates under detailed guidelines and is supported by multiple examples of human-authored questions, including stems, options, answers, and explanations, to ensure adherence to question design standards. Finally, the image matcher pairs images with questions by leveraging either direct image-question matching or matching based on image descriptions, depending on the available data. This Multi-Agent approach enables large-scale generation of high-quality question datasets efficiently and with minimal human intervention.
2.3 Validation, Stage \@slowromancapiii@
This subsection outlines the stringent validation procedures implemented to ensure the quality and relevance of the data used in model evaluation. Two critical types of validation are performed: Agent Validation and Expert Validation.
Agent Validation.
This stage aims to eliminate duplicate questions and verify the strength of the correlation between the visual content and the corresponding text. First, the Validation Agent calculates the textual similarity between questions and filters out highly redundant ones based on predefined criteria to ensure the diversity of the dataset. Next, for all constructed data, it is necessary to ensure that 1) without providing an image, the question cannot be correctly answered based solely on the text, and 2) without providing a question, the image alone cannot lead to the correct answer. This requirement stems from our aim to scrutinize the model’s capabilities from a multimodal perspective. If a single modality, text or images, suffices to answer a question, it compromises the quality of the dataset and its ability to comprehensively evaluate the performance of MLLMs. To meet these two requirements, the Validation Agent assesses the extent to which the question depends on the image. If the image is deemed unnecessary, the question is sent back to the Stage \@slowromancapii@ for revision; if it still fails to meet the requirements after multiple iterations, the question is discarded.
Expert Validation.
This stage focuses on expert evaluation of the data. For data annotated by experts, each entry must be 1) validated by other domain experts to confirm its appropriateness and lack of ambiguity, and 2) confirmed by all experts to ensure that it is free from sensitive issues. In the case of data generated by models, a rigorous evaluation by experts specializing in data generation is required to verify its accuracy and absence of ambiguity.
Furthermore, it requires the collective consent of all experts to confirm that the data are free of sensitive issues.
2.4 Implementation Details
Figure 4: Overview of HSSBench.
In our VGP, we leverage GPT-4o and GPT-4.1 due to their outstanding capabilities. Throughout all stages of LLM, we employed the Chain-of-Thought prompting strategy. For more detailed information on VGP, please refer to Appendix A.
2.5 Data Statistics
Our dataset comprises a total of 13,152 questions spanning six categories and 45 types. It supports evaluation in six different languages. The statistical data of HSSBench is illustrated in Figure 4, with a detailed description of the dataset provided in Appendix B.
3 Experiment
3.1 Experimental Setup
Models.
For open-source models, we selected Qwen2.5-VL-3B/7B/32B/72B Qwen2.5-VL , Qwen2-VL-72B, QVQ-72B-Preview Qwen2-VL , Deepseek-VL2-Tiny wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels , MiniCPM-o-2.6 yao2024minicpm , mPLUG-Owl3-2B/7B ye2024mplugowl3longimagesequenceunderstanding , Llava-onevision-7B li2024llavaonevisioneasyvisualtask , Llama3-llava-next-8b li2024llavanext-strong , InternVL3-8B zhu2025internvl3exploringadvancedtraining , InternVL2.5-8B-MPO chen2024expanding , Phi-3.5-Vision-Instruct abdin2024phi3technicalreporthighly , Janus-Pro chen2025janusprounifiedmultimodalunderstanding , Llava1.5 liu2024improvedbaselinesvisualinstruction . For closed-source commercial models, we utilized GPT-4o/4.1/4.1-mini/4.1-nano openai2024gpt4ocard .
Evaluation Settings.
1) We conducted evaluations using two types of prompts to examine the MLLMs performance under different prompting strategies. One approach involved prompting the model to directly output the answer without any intermediate reasoning, while the other required the model to generate a chain of thought before providing the final answer. 2) Additionally, we tested each question in six different language versions to investigate how the MLLMs’ performance varies across languages for the same question. 3) Furthermore, we optimized the question formats by designing two types of prompts for each item: multiple-choice and open-ended questions. This allowed us to assess whether the model can correctly answer questions without any explicit hints. 4) For the HSSBench evaluation, we employed two assessment methods and selected GPT-4o as a representative closed-source model and Qwen2.5-7B as a representative open-source model.A detailed comparative analysis of both models is provided in Appendix C.5.
3.2 Results
Overall Results.
Table 1 presents the overall results of various open-source and closed-source MLLMs evaluated on HSSBench in the context of English language tests. Ct. and Dt. represent two types of prompts: Chain-of-Thought (COT) prompts and direct response prompts. C. and O. denote question types: multiple-choice and open-ended questions. "Human" refers to the average final scores achieved by experts in various fields. Detailed experimental results for Dt. and are presented in Appendix C.1.
All of our initial generated data consisted of multiple-choice questions. Since some questions cannot be reasonably converted to open-ended formats, changing all of them into open-ended questions would lead to poor-quality data. This approach would be unfair for evaluating the model and would fail to accurately reflect differences in model performance. Therefore, we chose to modify only those questions that can still be answered meaningfully as open-ended questions to serve as the evaluation data for open-ended question performance.
The experimental results highlight the performance differences between the various models. Of the open-source models, Qwen2.5-VL-72B-Instruct delivers the highest performance, although it still falls short of surpassing closed-source models in open-ended question tests. The GPT-4.1 series achieves state-of-the-art performance in most tasks, with a particularly impressive accuracy of 39.97% in open questions, almost double that of other closed-source models. In contrast, some open-source models show considerably lower performance.
Although data in the HSS domain was generally considered common knowledge, our study demonstrates that from a multimodal perspective, these models require more diverse training data and targeted fine-tuning to effectively address related tasks. We also observe that model performance varies by language. Multilingual evaluation results are provided in the Appendix C.3.
Model Geography Economy Culture Social Sciences History Art All
Ct.C. Ct.O. Ct.C. Ct.O. Ct.C. Ct.O. Ct.C. Ct.O. Ct.C. Ct.O. Ct.C. Ct.O. Dr.C. Dr.O. Ct.C. Ct.O.
Random 24.93 0.00 21.92 0.00 25.00 0.00 24.90 0.00 24.91 0.00 25.00 0.00 24.62 0.00 24.62 0.00
Human 94.14 - 93.06 - 92.99 - 94.44 - 93.84 - 95.53 - 93.83 - 93.83 -
Open-source LLM (Scale <<< 10B)
Qwen2.5-VL-3B-Instruct 35.25 11.74 28.56 16.36 34.61 1.85 39.46 11.50 36.23 11.07 34.55 3.62 29.01 9.94 34.99 9.33
Qwen2.5-VL-7B-Instruct 40.69 21.60 41.19 31.31 30.19 4.63 42.86 19.00 40.62 19.67 33.42 11.31 37.88 15.21 38.19 17.89
Llava-onevision-7b 32.05 7.04 32.23 11.68 31.02 1.85 36.63 3.00 27.07 4.51 32.92 3.17 36.20 5.73 31.56 5.20
Llama3-llava-next-8b 27.59 4.23 19.82 6.54 30.53 3.24 32.89 6.50 26.87 8.61 29.26 5.43 31.20 6.50 27.93 5.81
InternVL3-8B 42.12 10.80 33.70 16.36 38.69 7.41 48.30 13.00 45.09 12.70 38.61 13.57 42.14 12.27 41.42 12.31
InternVL2.5-8B-MPO 37.24 17.37 34.07 20.09 35.00 8.33 43.35 15.50 40.54 16.39 35.99 13.57 39.30 11.77 37.68 15.21
Phi-3.5-Vision-Instruct 25.55 9.39 26.80 13.55 29.01 3.70 28.78 3.00 20.31 7.79 28.81 4.07 35.89 10.32 26.04 6.96
Janus-Pro 29.11 8.45 22.54 6.07 41.00 6.02 34.65 12.00 29.75 10.25 33.22 7.69 30.03 8.49 31.66 8.41
Deepseek-VL2-Tiny 5.78 3.29 3.77 4.95 14.12 0.00 6.35 3.00 6.37 2.46 13.03 16.74 29.86 3.42 8.23 5.09
mPLUG-Owl3-2B 25.83 4.69 24.63 1.35 31.83 3.69 29.99 3.00 25.96 4.51 29.43 4.07 28.73 4.02 27.71 3.57
mPLUG-Owl3-7B 30.60 7.04 33.15 9.91 13.73 2.30 36.40 7.00 27.64 4.10 25.69 2.30 33.01 6.68 27.52 6.23
MiniCPM-o-2.6 26.02 5.98 19.08 7.22 22.22 3.57 26.83 8.59 25.96 4.03 22.25 5.38 3.70 5.71 24.11 5.71
Llava1.5 12.75 4.03 7.38 2.68 10.74 3.36 7.38 5.37 10.74 4.03 10.74 6.04 8.06 4.03 9.96 4.25
Open-source LLM (Scale >>> 10B)
Qwen2.5-VL-32B-Instruct 52.48 21.33 52.79 6.67 38.94 8.00 53.87 24.00 57.03 26.67 39.20 3.33 48.38 15.89 50.75 15.00
Qwen2-VL-72B-Instruct 50.74 17.86 52.55 32.65 45.91 8.62 50.19 16.07 55.53 16.36 41.34 16.36 54.22 20.43 49.39 17.21
Qwen2.5-VL-72B-Instruct 55.59 13.33 53.83 37.33 41.49 7.33 57.77 17.57 60.30 28.19 40.84 14.67 54.17 18.17 51.87 19.73
QVQ-72B-Preview 19.93 3.33 21.87 17.33 29.67 3.33 26.54 7.43 28.86 18.79 23.80 9.33 25.60 10.37 24.69 9.92
Closed-source LLM
GPT-4o 46.88 22.07 52.97 35.14 45.61 14.29 45.26 16.00 48.36 15.98 43.42 12.67 46.09 20.05 46.88 19.36
GPT-4.1 39.81 40.38 48.08 52.70 48.95 24.88 35.36 49.25 41.91 48.77 43.51 23.53 45.02 25.38 42.66 39.97
GPT-4.1-mini 47.67 34.27 58.27 49.10 48.26 20.74 45.05 36.68 47.84 36.89 43.88 23.53 45.75 24.32 48.03 33.59
GPT-4.1-nano 33.21 30.74 39.71 41.44 38.70 7.83 37.52 28.14 34.01 30.74 35.83 20.36 36.33 21.12 35.83 26.22
Table 1: Scores (%) of MLLMs on HSSBench (EN-\@slowromancapi@). The highest and second highest scores are marked in blue and green, respectively.
Model performance on different categories.
Among the six categories evaluated on HSSBench (EN-\@slowromancapi@), the economic-related tasks consistently emerge as the most challenging for the models. The average score of all models in this category is the lowest, indicating that addressing economic problems requires a deep understanding of various economic theories and the ability to apply them in complex reasoning. However, current open-source MLLMs still exhibit significant deficiencies in these aspects. In contrast, closed-source models perform exceptionally well in economically related tasks, and this advantage may stem from their exposure to large amounts of high-quality training data in the economic domain.
On the other hand, the Geography category appears to be the easiest task for the models, with the highest average scores observed across the board. This trend implies that geographic knowledge, which is often more factual and less abstract compared to other humanities and social sciences domains, is better captured by the training data and reasoning capabilities of the models.
Interestingly, in certain categories of multiple-choice tasks, such as Culture and Social Sciences, some open-source models outperform their closed-source counterparts. For example, larger open-source models like Qwen2.5-VL-32B and Qwen2.5-VL-72B demonstrate competitive or even superior results compared to closed-source models like GPT-4o in these domains. This phenomenon may be partially attributed to the fact that most of the data experts we invited are Chinese. Although the evaluation was conducted in English, it is possible that Qwen benefits from a training data advantage related to Chinese content, which could have influenced its superior performance in these tasks. It also indicates that the gap between open-source and closed-source models is narrowing in specific HSS tasks.
3.3 Qualitative Analysis
Comparison between Direct Answer and COT Prompting.
Table 1 demonstrates that the performance of various MLLMs on HSSBench using direct answer prompts and COT prompts exhibits subtle differences. COT prompts do not always yield positive effects, and certain models perform better with direct answer prompts. This suggests that extending the reasoning process might mislead the models.
In HSS tasks, the COT method exhibits limitations for certain models. It not only fails to deepen the models’ understanding but may also cause them to focus on irrelevant information, thereby interfering with attention to critical knowledge. Specifically, COT prompts exacerbate hallucination issues in certain models, where reasoning flaws in textual analysis and misinterpretation of visual inputs lead to the generation of incorrect background knowledge during step-by-step analysis. This causes reasoning to deviate from the correct answer. For example, when tackling geographic questions, models often struggle to accurately interpret spatial elements such as location markers and contour lines—visual features primarily composed of points and lines—resulting in analytical outcomes that contradict the actual image content.
Even when notable errors do not occur in intermediate reasoning steps, the final summarization phase can suffer due to excessive information generation, which exceeds the model’s ability to effectively weigh the importance of answer options, leading to prediction mistakes. This highlights the importance of integrating reasoning steps cohesively. Further detailed analysis in Appendix D reveals that many models fail to internalize visual knowledge during the divergent thinking process of HSS tasks.
Comparison between Multiple-choices and Open-ended.
We selected questions from the dataset that can be reformulated as open-ended to evaluate model performance without the guidance of provided answer options. The results indicate that HSS tasks pose significant challenges for current MLLMs, with only a few models achieving accuracy above 15% in the open-ended setting. This result aligns with our expectations: even experts find it difficult to provide the correct answer without predefined options and background knowledge.
The presence of answer options offers models valuable prior knowledge, effectively constraining the scope of possible answers. In contrast, open-ended questions remove this constraint, causing the model’s reasoning to become highly divergent and often leading to answers that deviate substantially from the ground truth. This suggests that models primarily rely on shallow visual attributes, such as object size, type, or motion, while overlooking deeper symbolic information embedded in cultural context, academic theories, or spatiotemporal cues.
These observations highlight two key limitations of current MLLMs: first, a lack of comprehensive understanding and memory of real-world knowledge; second, difficulty in effectively integrating visual features with textual information. Moreover, the performance degradation observed when introducing chain-of-thought prompting further underscores the models’ struggles with internalizing and retrieving complex knowledge.
The role of accurate answers.
In multiple-choice settings, the presence of answer options inevitably provides additional cues that can influence model performance. To further investigate whether models truly understand the questions rather than relying on superficial cues, we introduced confusing options, specifically, the choice "None of the above," while keeping the correct answer unchanged. This experiment was conducted on a subset of 150 samples randomly selected from each category. The results are illustrated in Table 5.
For lower-performing models, the introduction of confusing options led to a noticeable decline in performance. Upon closer examination of the models’ reasoning processes after adding these distractors, two main patterns emerged. In some cases, models were directly swayed by the confusing option and selected "None of the above" incorrectly. More intriguingly, in other cases, although the models did not explicitly choose the confusing option, their reasoning became less coherent.
This behavior indicates that current models struggle to accurately assess the relative credibility of different answer choices and to make reliable judgments when faced with conflicting information. Consequently, their performance deteriorates in the presence of carefully designed distractors.
Visual Information Extraction.
To investigate the impact of visual information loss on the performance of MLLMs, we designed two complementary experiments. In the first experiment, we used GPT-4.1 to generate detailed textual descriptions of the images. In the second experiment, we invited domain experts to produce comprehensive and precise annotations for each image. The experimental results are summarized in Table 6.
Specifically, we converted the content of images into textual descriptions and provided only these texts, along with the questions and answer options, to the models. The accuracy of most models declined compared to when the full images were available, indicating that direct visual input contains critical information that is partially lost in the text conversion process.
However, considering that even when models receive the full images, they may not fully comprehend certain fine-grained visual details, we further involved domain experts to produce detailed annotations of the images. These expert annotations distilled key visual information into precise textual descriptions. When provided with these expert-annotated texts instead of the images, the models’ accuracy notably improved relative to the plain text conversion scenario, despite the absence of the original images.
Interestingly, some models even surpassed their previous upper-bound performance in domains such as Culture and Social Sciences. This improvement suggests that expert annotations effectively guide the models to focus on critical visual cues by translating them into semantically rich and relevant textual information, thereby facilitating stronger connections with domain knowledge. These findings corroborate our earlier observations on open-ended questions, highlighting inherent limitations in current MLLMs’ abilities to retrieve and interpret visual information comprehensively.
4 Limitations
Our HSSBench is designed to evaluate MLLMs’ capabilities through HSS reasoning tasks. The type within our dataset reflects our effort to collect data from as many perspectives as possible. However, we acknowledge that the scope of knowledge covered is inherently limited by the contributors’ expertise and cannot encompass all aspects of HSS. Moreover, HSS currently lacks a universally accepted definition. Our evaluation is based on six major categories that, while comprehensive, do not cover the full spectrum of HSS-related fields.
In addition, some data points involve recent social events and the latest economic knowledge. We recognize that such data are time sensitive. Although the answers are accurate at present, future developments may lead to different interpretations. For these items, we impose temporal and spatial constraints to ensure their correctness within the current context.
Although these limitations are relevant, we believe that they extend beyond the scope and original intent of our dataset. We hope that future work will address these challenges in more detail.
5 Related Work
In recent years, significant progress has been made in the development of multimodal benchmarks and methodologies. Numerous datasets have emerged that assess models from various perspectives, which can be broadly categorized into Generation Benchmarks liu2024mmbench ; liu2023visualinstructiontuning ; meng2024vgavisionguiassistant ; han2024instinctivebiasspuriousimages ; qian2024easyfoolmultimodalllms ; xu2023lvlmehubcomprehensiveevaluationbenchmark ; yin2023lammlanguageassistedmultimodalinstructiontuning ; zeng2023matterstraininggpt4stylelanguage ; wu2024qbenchbenchmarkgeneralpurposefoundation ; luo-etal-2024-codis , Reasoning Benchmarks wang2024measuringmultimodalmathematicalreasoning ; lu2024mathvistaevaluatingmathematicalreasoning ; li2023scigraphqalargescalesyntheticmultiturn ; liang2024scemqascientificcollegeentrance ; song2025csbench , and Application Benchmarks fan2022minedojobuildingopenendedembodied ; rawles2023androidwildlargescaledataset ; chen2023endtoendembodieddecisionmaking ; sermanet2023robovqamultimodallonghorizonreasoning ; chen2024egoplanbenchbenchmarkingmultimodallarge ; koh2024visualwebarenaevaluatingmultimodalagents ; you2024ferretuigroundedmobileui . Generation benchmarks cover a wide range of tasks, with the core objective of comprehensively evaluating the multi-dimensional performance of MLLMs. They not only include reasoning challenges but also address areas related to humanities and social sciences. In contrast, reasoning benchmarks are designed primarily around mathematical and scientific problems, seeking to rigorously assess the capabilities of MLLM in logical and analytical reasoning. Among the various evaluation formats, multiple-choice datasets are the most prevalent, owing to their simplicity in evaluation and ease of comparison zeng2023matterstraininggpt4stylelanguage .
The proliferation of these data sets has significantly accelerated research progress, serving both as training resources and as tools for assessing the multifaceted competencies of MLLMs. Advances in data set construction are particularly exciting, ranging from human-annotated high-quality datasets romero2024cvqaculturallydiversemultilingualvisual to those generated through pipelines built in LLM chandrasegaran2024hourvideo1hourvideolanguageunderstanding . Despite these achievements, several critical limitations remain. 1) Most multimodal datasets cover a wide range of categories, but suffer from issues such as limited data sources, relatively simple questions, and insufficient image information. 2) Many large-scale datasets are collected through web crawling without thorough manual annotation and verification, which can introduce biases in evaluation results. 3) Current reasoning datasets predominantly focus on STEM tasks, relying heavily on scientific and mathematical data to assess reasoning capabilities. In contrast, the benchmark we propose is created through interactions among expert agents, targeting the HSS domain. In contrast, the benchmark we propose focuses on the HSS domain, emphasizing tasks that analyze and understand the abstract concepts embedded in images. This approach fills a notable gap in the field and lays the foundation for advancing model performance in this underexplored area.
6 Conclusion
We present HSSBench, a novel benchmark dataset constructed through a multi-agent pipeline involving experts from diverse fields. HSSBench is designed to rigorously evaluate the true mastery of tasks by models within the HSS domain. The dataset comprises six categories, each derived from raw data collected from repositories in the six official languages of the United Nations and subsequently processed to generate task-specific data.
We then carried out comprehensive benchmarking of various MLLMs using HSSBench. Our results reveal that HSSBench poses significant challenges to all tested models, which exhibit poor performance on reasoning tasks in the HSS domain. In particular, the accuracy of the model decreases substantially when answer choices are not provided as prompts. We hope that releasing HSSBench will encourage the AI community to place greater emphasis on reasoning over non-STEM data, thereby advancing research on MLLMs from this important perspective.
Appendix A More VGP Details
A.1 Annotation Platform
Figure 5: The interface for data construction and validation, allowing experts to use a visual interface to assist their work.
We use a simple visual interface as our annotation platform. For question input, experts can upload or write questions in the form. The interface is shown in Figure 5. During the validation process, experts can see all the data submitted by other experts. They can select entries to view detailed previews of the entries.
Appendix B Dataset Details
This section provides detailed information about our benchmark designed to evaluate MLLMs’ visual comprehension abilities through multiple-choice questions. The benchmark spans six major categories: Economy, Art, Culture, Social sciences, History, and Geography, each containing various specialized subtypes.
B.1 Human Performance
To establish a performance baseline, we asked three relevant experts in each domain to spend a significant amount of time answering the entire dataset. Table 2 summarizes their performance across the six categories.
Category Expert 1 Expert 2 Expert 3 Overall
Economy 96.02 91.70 94.70 94.14
Art 93.12 92.61 93.44 93.06
Culture 92.50 90.83 95.64 92.99
Social sciences 94.49 91.77 97.07 94.44
History 95.91 91.86 93.75 93.84
Geography 94.99 92.78 95.81 94.53
Average 94.84 91.76 95.40 93.83
Table 2: Scores (%) of Experts.
Human experts demonstrated high proficiency across all categories, with overall accuracy ranging from 92.99% to 94.53%. The highest individual performance was observed in Social sciences by Expert 3 (97.07%), while the lowest was in Culture by Expert 2 (90.83%). This high level of human performance establishes a challenging benchmark for evaluating large language models.
B.2 The details of types
Our dataset is organized into six major categories with various subtypes in each category. Table 3 presents the detailed breakdown of the dataset structure and the count of questions in each type. We allow each data entry to have multiple types because the intersection of knowledge across disciplines is essential.
The dataset exhibits varying distributions across categories, with Geography containing the largest number of questions and Economy the smallest. Within categories, there are also significant variations in subtype representation, reflecting the natural distribution of content within these domains.
Category Type Count
Economy Microeconomics 1,193
Macroeconomics 163
Labor Economics 105
Environmental Economics 63
International Trade 63
Resource Economics 44
International Finance 32
Art Folk Art 1,051
Eastern Art 1,041
Western Art 672
Spiritual Art 630
Applied Art 452
Entertainment Art 298
Elite Art 303
Culture Regional Culture 1,790
Material Culture 1,382
Intangible Culture 1,218
Indigenous Culture 733
Institutional Culture 266
World Culture 228
Social sciences Sociology 950
Political Science 550
Psychology 339
Anthropology 190
Economics 147
Education 135
Philosophy 127
Law 99
Ethics 61
Journalism 24
History 18
History Asian History 1,586
World Modern History 885
World Ancient History 754
European History 602
World Contemporary History 346
American History 295
African History 147
Oceanian History 82
Geography Physical Geography 2,361
Regional Geography 1,694
Descriptive Geography 1,682
Explanatory Geography 1,348
Human Geography 1,319
Predictive Geography 75
Table 3: Dataset Taxonomy and Question Distribution.
B.3 Most-Frequent Words in the Questions
We analyzed the most frequent content words in the questions in all categories to understand the linguistic characteristics and tasksur dataset. Figure 6 shows the word clouds for the most frequent words in HSSBench per category, excluding common stop words. The word frequency analysis reveals distinct patterns across categories:
1) Visual observation terms dominate in Art and Culture categories ("observe", "picture", "shown", "scene"), indicating a focus on visual analysis tasks.
2) Economy questions frequently use technical terms ("price", "firm", "cost", "demand", "market", "marginal"), reflecting domain-specific concepts.
3) Geography questions heavily employ spatial and diverse visual comprehension capabilities ("diagram", "map", "area", "distribution"), emphasizing spatial reasoning.
(a) Art
(b) Culture
(c) Economy
(d) Geography
(e) History
(f) Social sciences
Figure 6: Word Cloud in HSSBench per category.
Appendix C Experimental details
C.1 Details of Results
Table 4 provides additional details on the experimental results presented in Table 1. The table below reports the overall accuracy of the models under the DIRECT prompt and presents a detailed breakdown of the performance for each category under the same prompt.
Model Geography Economics Culture Social Sciences History Art All
Dr.C. Dr.O. Dr.C. Dr.O. Dr.C. Dr.O. Dr.C. Dr.O. Dr.C. Dr.O. Dr.C. Dr.O. Dr.C. Dr.O. Ct.C. Ct.O.
Random 24.93 0.00 21.92 0.00 25.00 0.00 24.90 0.00 24.91 0.00 25.00 0.00 24.62 0.00 24.62 0.00
Human 94.14 - 93.06 - 92.99 - 94.44 - 93.84 - 95.53 - 93.83 - 93.83 -
Open-source LLM (Scale <<< 10B)
Qwen2.5-VL-3B 29.84 11.27 29.44 17.76 27.83 5.09 35.22 10.00 26.91 8.61 26.78 7.24 29.01 9.94 34.99 9.33
Qwen2.5-VL-7B 40.07 19.25 38.18 30.37 30.29 7.87 44.91 11.00 41.20 14.75 32.33 8.14 37.88 15.21 38.19 17.89
Llava-onevision-7b 37.46 6.10 31.20 10.75 38.64 1.39 40.95 7.00 31.65 4.51 37.33 4.98 36.20 5.73 31.56 5.20
Llama3-llava-next-8b 29.98 6.10 21.81 5.61 35.69 6.48 35.36 6.00 29.48 9.02 34.46 5.43 31.20 6.50 27.93 5.81
InternVL3-8B 42.88 13.62 37.08 17.97 37.71 6.91 48.94 12.00 47.22 14.34 37.48 8.60 42.14 12.27 41.42 12.31
InternVL2.5-8B-MPO 38.08 14.08 34.73 16.36 39.13 6.94 45.69 10.50 40.82 11.07 38.27 11.76 39.30 11.77 37.68 15.21
Phi-3.5-Vision-Instruct 29.84 10.33 29.44 11.21 27.83 10.65 35.22 7.00 26.91 11.48 26.78 10.86 29.01 10.32 26.04 6.96
Janus-Pro 27.57 8.92 16.01 7.01 42.38 6.02 31.26 9.00 28.62 12.30 32.33 7.24 30.03 8.49 31.66 8.41
Deepseek-VL2-tiny 25.44 5.16 17.38 3.15 40.17 0.46 30.33 2.50 29.97 4.51 35.19 4.52 29.86 3.42 8.23 5.09
mPLUG-Owl3-2B 25.27 4.23 30.48 2.70 32.56 2.30 34.17 4.50 26.88 6.15 30.48 4.07 28.73 4.02 27.71 3.57
mPLUG-Owl3-7B 32.71 7.04 33.08 6.31 33.45 4.61 37.31 5.50 29.97 8.20 33.73 8.14 33.01 6.68 27.52 6.23
MiniCPM-o-2.6 3.48 1.67 3.90 14.16 2.06 3.70 5.75 3.03 4.72 2.27 3.02 8.70 3.70 5.71 24.11 5.71
Llava1.5 3.36 2.68 10.07 5.37 3.36 2.01 8.05 4.70 8.05 1.34 15.44 8.05 8.06 4.03 9.96 4.25
Open-source LLM (Scale >>> 10B)
Qwen2.5-VL-32B 52.54 16.00 39.38 8.00 39.87 10.00 53.09 26.67 57.52 24.67 44.92 10.00 48.38 15.89 50.75 15.00
Qwen2.5-VL-72B 56.71 16.00 51.67 32.00 57.77 5.33 47.43 16.89 59.26 24.16 63.12 14.67 54.17 18.17 51.87 19.73
QVQ-72B-Preview 21.53 3.33 18.38 22.00 29.57 2.67 26.26 10.14 29.58 12.75 23.85 11.33 25.60 10.37 24.69 9.92
Qwen2-VL-72B 54.22 19.30 41.06 24.07 55.33 5.66 58.85 12.96 62.17 32.08 50.00 18.87 54.22 17.79 49.39 18.82
Closed-source LLM
GPT-4o 45.68 23.47 48.92 35.14 48.16 15.67 46.72 18.00 47.24 14.75 46.66 13.57 46.99 20.05 46.88 19.36
GPT-4.1 43.80 26.29 52.34 42.34 52.77 18.89 40.86 26.13 38.86 22.13 44.74 16.74 45.02 25.38 42.66 39.97
GPT-4.1-mini 45.14 26.29 54.92 45.95 52.08 15.21 41.84 22.11 39.34 18.85 44.70 17.65 45.75 24.32 48.03 33.59
GPT-4.1-nano 33.18 20.66 34.96 36.04 45.61 14.75 38.42 15.58 32.69 20.49 36.47 18.55 36.33 21.12 35.83 26.22
Table 4: Scores (%) of MLLMs on HSSBench (EN-\@slowromancapii@).
(a) Effect under multiple-choice question setting
(b) Effect under open-ended question setting
Figure 7: Partial MLLMs’ results under CoT prompt settings.
C.2 Details of Contrast Experiment
Table 5 and Figure 8 provides a detailed presentation of the experimental results after adding a confounding option. We sampled 900 data points. "Conf." indicates that, in addition to the given options, an extra option—"None of the above answers is correct"—was added. The model’s output performance was then compared under these conditions.
Model Geography Art Culture Social Sciences History Economy Total
Ct.C. Ct.C.Conf. Ct.C. Ct.C.Conf. Ct.C. Ct.C.Conf. Ct.C. Ct.C.Conf. Ct.C. Ct.C.Conf. Ct.C. Ct.C.Conf. Ct.C. Ct.C.Conf.
Qwen2.5-VL-7B 48.00 44.67 32.12 31.69 29.33 26.67 39.86 34.00 55.33 49.33 35.51 36.05 40.21 37.12
InternVL3-8B 50.67 46.00 41.55 34.01 45.33 49.33 53.38 48.67 60.00 56.00 39.86 45.99 49.37 46.01
MiniCPM-o 27.51 21.47 39.33 34.67 49.33 50.67 35.81 29.73 38.26 26.85 27.33 23.33 36.34 31.22
Qwen2.5-VL-32 50.83 49.83 46.10 40.90 41.55 30.99 59.61 51.92 61.98 56.77 50.94 53.77 51.85 47.74
QvQ-72B-Preview 16.67 16.67 29.33 28.00 35.78 29.67 26.35 18.91 28.18 27.51 25.33 24.66 26.98 24.41
Qwen2.5-VL-72 60.67 55.33 45.33 42.05 49.33 44.67 56.76 55.40 65.77 67.11 54.67 55.33 55.41 53.29
GPT-4.1-mini 46.67 47.33 56.67 54.67 52.67 47.33 40.00 42.67 54.67 44.00 45.33 42.00 49.33 46.33
Table 5: Contrast Scores (%) with confounding option of MLLMs on HSSBench.
(a) Art
(b) Culture
(c) Economy
(d) Geography
(e) History
(f) Social sciences
Figure 8: Contrast Scores (%) with confounding option of MLLMs on HSSBench.
Table 6 and Figure 9 presents the extraction of image information into text (where "De." indicates a detailed description of the image generated directly by GPT-4.1, and "De-H." indicates a detailed explanation provided by a domain expert for each image). The table below shows the detailed results of comparative experiments in which only the extracted image information, rather than the images themselves, was provided to the model.
Model Geography Art Culture Social Sciences History Economy
Ct.C. De.C. De-H.C. Ct.C. De.C. De-H.C. Ct.C. De.C. De-H.C. Ct.C. De.C. De-H.C. Ct.C. De.C. De-H.C. Ct.C. De.C. De-H.C.
Qwen2.5-VL-7B 39.33 40.67 41.33 43.48 37.41 40.14 30.67 30.00 36.00 37.16 39.33 42.67 44.00 46.00 51.33 35.77 35.21 38.73
InternVL3-8B 42.00 43.33 43.33 33.33 34.69 36.05 42.00 32.00 42.00 47.30 41.33 45.33 48.00 48.00 50.00 40.88 35.92 39.44
Qwen2.5-VL-32B-Instruct 46.00 44.67 52.00 35.33 44.00 43.33 42.67 34.00 47.33 48.67 44.67 52.67 44.00 52.67 62.67 26.67 52.00 55.33
QvQ-72B-Preview 31.33 36.67 42.00 26.67 30.67 38.00 28.67 27.33 36.00 34.67 34.00 36.00 29.33 40.67 38.67 19.33 37.33 33.33
Qwen2.5-VL-72B-Instruct 54.67 54.67 58.67 42.00 48.67 41.33 34.67 32.67 42.67 47.33 50.00 47.33 56.67 64.67 66.67 35.33 56.00 57.33
GPT-4.1-mini 52.00 50.00 56.00 43.33 42.67 52.67 50.00 40.67 54.00 42.67 40.00 48.67 48.67 46.67 54.67 55.33 53.33 60.00
Table 6: Contrast Scores (%) about Visual Information Extraction of MLLMs on HSSBench.
(a) Art
(b) Culture
(c) Economy
(d) Geography
(e) History
(f) Social sciences
Figure 9: Contrast Scores (%) about Visual Information Extraction of MLLMs on HSSBench.
C.3 Results in different languages
Table 1 and Table 4 present the test results on English-language data. To analyze the impact of different languages on the model, we also conducted large-scale experiments using Chinese-language data. The final experimental results are shown in Table 7 and Table 8 below.
Model Geography Economics Culture Social Sciences History Art All
Ct.C. Ct.O. Ct.C. Ct.O. Ct.C. Ct.O. Ct.C. Ct.O. Ct.C. Ct.O. Ct.C. Ct.O. Dr.C. Dr.O. Ct.C. Ct.O.
Random 24.93 0.00 21.92 0.00 25.00 0.00 24.90 0.00 24.91 0.00 25.00 0.00 24.62 0.00 24.62 0.00
Human 94.14 - 93.06 - 92.99 - 94.44 - 93.84 - 95.53 - 93.83 - 93.83 -
Open-source LLM (Scale <<< 10B)
Qwen2.5-VL-3B 40.49 10.80 28.89 19.81 36.30 0.93 42.43 9.50 43.91 11.89 33.85 4.98 35.45 11.79 35.45 9.65
Qwen2.5-VL-7B 48.43 15.96 35.00 25.00 27.60 2.78 49.36 19.50 50.54 15.98 32.90 6.79 43.12 13.17 41.86 14.24
Llava-onevision-7b 39.68 4.69 30.80 10.38 32.96 1.85 39.46 7.00 32.97 4.51 35.03 2.71 37.89 4.98 35.62 5.13
Llama3-llava-next-8b 27.29 3.29 19.82 3.77 33.55 0.93 32.89 2.50 26.84 3.28 29.48 2.26 31.18 5.21 28.35 2.68
InternVL3-8B 47.45 11.74 35.37 17.92 38.21 9.26 53.47 14.00 50.97 18.44 39.94 10.41 44.81 14.70 44.92 13.71
InternVL2.5-8B-MPO 41.14 11.74 36.11 23.58 36.20 6.94 48.09 16.50 47.63 6.94 38.35 10.86 42.72 14.01 41.45 14.55
Phi-3.5-vision-instruct 22.29 6.10 25.64 4.25 30.16 2.31 23.55 1.00 23.35 4.92 28.30 1.81 34.08 6.74 25.35 3.45
Janus-Pro 28.10 11.27 21.67 8.49 44.16 6.48 34.79 8.50 29.13 14.34 40.49 14.48 33.50 11.94 32.81 10.72
mPLUG-Owl3-2B 27.20 1.88 29.43 0.90 33.74 2.30 29.78 3.00 26.76 2.87 29.43 4.52 32.25 1.90 28.25 2.58
mPLUG-Owl3-7B 35.01 5.63 34.23 3.15 26.39 1.84 39.26 2.50 26.08 2.46 34.23 5.43 34.22 4.71 31.77 3.49
MiniCPM-o-2.6 26.97 7.69 19.50 4.88 35.21 6.25 31.23 3.90 30.27 6.76 30.20 5.98 1.29 6.83 29.04 5.94
Open-source LLM (Scale >>> 10B)
Qwen2.5-VL-32B 58.45 19.33 49.86 7.33 40.90 5.33 56.49 22.67 64.33 24.67 41.62 14.67 51.60 14.22 52.86 15.67
Qwen2.5-VL-72B 61.65 12.67 52.02 33.33 43.60 10.67 61.04 14.19 68.08 18.12 43.76 16.00 60.04 21.96 55.94 17.51
QVQ-72B-Preview 19.82 8.00 21.10 18.00 30.80 3.33 25.41 6.76 27.85 10.74 25.90 6.67 22.95 9.14 24.82 8.92
Qwen2-VL-72B 21.28 19.30 38.25 24.07 21.25 5.66 27.79 12.96 25.69 32.08 27.65 18.87 59.11 17.79 25.59 18.82
Closed-source LLM
GPT-4.1mini 47.53 35.21 55.55 55.41 44.43 18.43 44.07 34.17 47.92 43.03 42.50 30.32 44.19 24.54 46.78 36.32
Table 7: Scores (%) of MLLMs on HSSBench (ZH-\@slowromancapi@).
Model Geography Economics Culture Social Sciences History Art All
Dr.C.g Dr.O.g Dr.C.g Dr.O.g Dr.C.g Dr.O. Dr.C.g Dr.O.g Dr.C.g Dr.O.g Dr.C.g Dr.O.g Dr.C.g Dr.O.g Ct.C.g Ct.O.g
Random 24.93 0.00 21.92 0.00 25.00 0.00 24.90 0.00 24.91 0.00 25.00 0.00 24.62 0.00 24.62 0.00
Human 94.14 - 93.06 - 92.99 - 94.44 - 93.84 - 95.53 - 93.83 - 93.83 -
Open-source LLM (Scale <<< 10B)
Qwen2.5-VL-3B 39.51 12.21 29.26 13.21 30.94 7.87 40.03 14.50 38.36 11.48 30.03 11.76 35.45 11.79 35.45 9.65
Qwen2.5-VL-7B 52.52 17.84 33.97 20.75 27.60 4.63 51.13 15.00 52.13 11.48 31.17 9.95 43.12 13.17 41.86 14.24
Llava-onevision-7b 42.20 6.57 30.21 7.55 36.59 1.39 42.93 5.00 35.61 4.10 36.12 5.43 37.89 4.98 35.62 5.13
Llama3-llava-next-8b 29.61 4.69 19.97 6.60 38.80 5.09 34.37 5.00 28.86 3.28 34.54 6.33 31.18 5.21 28.35 2.68
InternVL3-8B 46.44 18.78 34.27 20.75 40.08 6.94 52.90 18.00 51.05 14.34 40.19 9.95 44.81 14.70 44.92 13.71
InternVL2.5-8B-MPO 42.26 18.78 31.32 14.15 40.72 7.87 48.87 17.50 49.11 12.30 40.78 14.03 42.72 14.01 41.45 14.55
Phi-3.5-vision-instruct 29.92 6.57 27.49 7.08 45.58 7.87 32.89 4.50 29.71 6.56 40.68 7.69 34.08 6.74 25.35 3.45
Janus-Pro 27.29 9.86 18.50 12.74 50.83 14.35 32.18 8.50 29.44 10.25 43.21 15.84 33.50 11.94 32.81 10.72
mPLUG-Owl3-2B 31.25 0.94 23.73 2.71 35.95 0.46 37.17 3.00 31.89 2.46 33.18 2.71 32.25 1.90 28.25 2.58
mPLUG-Owl3-7B 35.85 7.51 35.65 6.79 35.70 1.38 40.59 3.00 28.17 4.10 35.65 6.79 34.22 4.71 31.77 3.49
MiniCPM-o-2.6 0.93 7.55 0.77 8.06 1.42 1.61 1.99 5.41 0.60 20.00 2.42 2.27 1.29 6.83 29.04 5.94
Open-source LLM (Scale >>> 10B)
Qwen2.5-VL-32B 56.29 20.67 42.83 6.00 41.83 4.00 57.42 18.00 63.00 25.33 42.21 11.33 51.60 14.22 52.86 15.67
Qwen2.5-VL-72B 68.15 22.67 48.40 30.00 50.07 12.00 63.73 20.27 71.02 31.54 49.02 15.33 60.04 21.96 55.94 17.51
QVQ-72B-Preview 19.40 6.67 17.90 20.00 29.72 4.00 24.34 11.49 28.13 8.05 23.44 4.67 22.95 9.14 24.82 8.92
Qwen2-VL-72B 64.42 28.07 41.83 16.67 55.98 7.55 62.00 14.81 68.77 22.64 51.49 16.98 59.11 17.79 25.59 18.82
Closed-source LLM
GPT-4.1mini 45.93 26.76 53.31 43.69 46.25 14.29 39.61 20.10 39.26 24.18 42.09 17.65 44.19 24.54 46.78 36.32
Table 8: Scores (%) of MLLMs on HSSBench (ZH-\@slowromancapii@).
Figure 10 illustrates the performance of various models under four different prompt configurations.
Figure 10: Comparison of Model Performances Across Four Prompt Settings.
Table 9 and Figure 11 shows the performance of the model on datasets in different languages. Our dataset includes data organized in six languages. The table below presents the experimental results on a stratified sample of 900 instances.
Model Arabic Chinese English French Russian Spanish
InternVL3-8B 35.20 49.37 42.38 39.37 38.70 39.21
Qwen2.5-VL-7B-Instruct 33.74 40.21 38.37 34.08 35.55 34.83
Qwen2.5-VL-32B-Instruct 46.94 54.79 50.24 47.57 48.51 50.24
Qwen2.5-VL-72B-Instruct 49.72 55.41 51.28 51.28 48.27 48.83
QVQ-72B-Preview 33.98 36.89 39.80 37.86 40.77 38.83
GPT-4.1-mini 41.33 44.89 41.67 41.78 41.44 46.89
Average 40.82 46.93 43.96 41.99 42.87 43.47
Table 9: Contrast Scores (%) of MLLMs on HSSBench for six UN languages and six models.
Figure 11: Contrast Scores (%) of MLLMs on HSSBench for six UN languages and six models.
C.4 Prompt for Model Inference
Table 10 details the configurations of the four prompts employed in our experiments, specifying the presence or absence of a CoT prompt and indicating whether the questions were open-ended or multiple-choice.
Prompt for Model Inference
w/ MC w/ CoT Question: [question] Options: [options] Think step by step to determine the correct answer. End your response with [[X]] where X is your final answer (A, B, C, D or E).
w/ MC w/o CoT Question: [question] Options: [options] Give the correct answer directly. End your response with [[X]] where X is your final answer (A, B, C, D or E).
w/o MC w/ CoT Question: [question] Think step by step to determine the correct answer. End your response with [[X]] where X is your final answer.
w/o MC w/o CoT Question: [question] Give the correct answer directly. End your response with [[X]] where X is your final answer.
Table 10: Prompt for model inference.
C.5 Evaluation for model’s output
Table 11 presents the evaluation prompts employed to assess the accuracy of the model’s responses.
Prompt for Model Evaluation
w/ MC You are an evaluation assistant. Please determine whether the answers output by the model below are correct. Question: [question] Options: [options] Correct answer: [correct answer] Model output content: [model output] Please extract its final answer from the model output and determine whether it is consistent with the content of the correct answer. If the answer is correct, reply with "1". Otherwise, reply with "0". Only reply with numbers and no superfluous content.
w/o MC You are an evaluation assistant. Please determine whether the answers output by the model below are correct. Question: [question] Correct answer: [correct answer] Model output content: [model output] Please extract its final answer from the model output and determine whether it is consistent with the content of the correct answer. Attention! If the meanings are roughly consistent, it can be regarded as correct. It is not necessary to be strictly and completely consistent. One needs to carefully consider whether the model’s response is reasonable and correct. If the answer is correct, reply with "1". Otherwise, reply with "0". Only reply with numbers and no superfluous content.
Table 11: Prompt for model evaluation.
In our evaluation process, we employed both expert sampling and automated assessment using GPT models. To verify the consistency of the evaluation results produced by the models, we randomly sampled 2,000 data points from the overall experiment, including 1,000 multiple-choice questions and 1,000 open-ended questions. These samples were independently evaluated by three human experts, as well as by Qwen2.5-7B and GPT-4.1.
For both multiple-choice and open-ended questions, the overall agreement among the two models and the three experts, measured by Fleiss’ Kappa, was 96.82% and 94.79%, respectively. The agreement between the two models alone, measured by Cohen’s Kappa, was 97.32% for multiple-choice questions and 94.79% for open-ended questions.
These results demonstrate that the consistency between the three experts and the two models remains at a reliable level. This indicates that either evaluation method yields results that are largely consistent with those reported in the paper, thereby ensuring the reproducibility of our findings.
Appendix D More Cases
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We have detailed in the abstract and introduction that HSS-related issues still pose significant challenges for large models. Our contribution lies in proposing a large-scale, high-quality benchmark to evaluate the performance of large models on HSS tasks.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In Section 4, we discuss two limitations of our work: 1) Since there is no fixed set of classification tasks in the HSS field, our dataset does not cover all possible tasks within HSS; 2) Some data in certain HSS subfields have spatiotemporal relevance, so we can only ensure that the questions in our dataset are correct at the present time.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: Our work is focused on datasets and does not involve theoretical assumptions or proofs.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Our experimental setup prompts are described in detail in Appendix C.4 and  C.5. We also verify the consistency of evaluation across different models; that is, similar results can be achieved using other open-source models as with our closed-source model tests. Our work can be reproduced on our dataset from both the testing and evaluation perspectives.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have open-sourced our dataset and implementation code on both GitHub and Hugging Face.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In Section 3.1, we provide a detailed description of our experimental design. Our experimental setup prompts are described in detail in Appendix C.4 and  C.5.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: In the Appendix C.5, we discuss the consistency among human evaluation, GPT evaluation, and Qwen evaluation. The consistency is within an acceptable range, demonstrating that the choice of evaluation method does not have a significant statistical impact on the final results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [N/A]
  39. Justification: Our benchmark work is not related to computing resources; as long as it can run, that’s sufficient.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our work complies with ethical standards, and all experts involved in data production are members of our team. Moreover, the images we use are either drawn by our experts or sourced from open-source data or websites. Our research does not have any harmful social impact.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Throughout this paper, we emphasize the positive social impact of our work. Our evaluation dataset can help the AI community assess the performance of large models on HSS tasks, while also highlighting the current insufficient attention given to HSS tasks. We believe our work can actively advance the development of MLLMs.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: The construction of our dataset involves the use of multiple agents for web data crawling. However, we have implemented numerous restrictions during the crawling process, and image and text information can only be obtained from websites that we have selected and approved. We assure you that there are no risks associated with the use of our networked multi-agent system.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We obtained data from https://www.unesco.org/zh/open-access/cc-nc-sa and comply with the CC BY-NC-SA license.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our dataset contains 13,152 image-text pairs, and each entry supports evaluation in the six official languages of the United Nations
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our dataset work does not involve crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our dataset work does not involve crowdsourcing or research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v2

@@
 1 Introduction
-Multimodal large language models (MLLMs) achiam2023gpt ; team2024gemini have demonstrated remarkable performance across a wide range of tasks, recently achieving or even exceeding human-level capabilities in many of them.
+Figure 2: An example of cross-modal knowledge transfer issues in MLLMs within the HSS domain.
+They struggle to associate "Business Penmanship" font knowledge with relevant images or recognize fonts in image text.
+Multimodal large language models (MLLMs) Achiam et al.
+(2023); Team et al.
+(2024) have demonstrated remarkable performance across a wide range of tasks, recently achieving or even exceeding human-level capabilities in many of them.
 As the performance of MLLMs continues to improve, conducting a comprehensive evaluation of their capabilities has become increasingly essential.
-In recent times, numerous benchmarks for evaluating MLLMs have emerged hendrycks2020measuring ; yue2024mmmu ; liu2024mmbench ; saikh2022scienceqa ; zhang2024cmmmu .
+In recent times, numerous benchmarks for evaluating MLLMs have emerged Hendrycks et al.
+(2020); Yue et al.
+(2024); Liu et al.
+(2024b); Saikh et al.
+(2022); Zhang et al.
+(2024).
 These tasks are designed to assess the models’ ability to jointly understand and reason across multiple modalities, such as images and text, from various perspectives.
-Specifically, most multimodal benchmarks are designed either from a general perspective liu2024mmbench ; liu2023visualinstructiontuning ; meng2024vgavisionguiassistant ; han2024instinctivebiasspuriousimages ; qian2024easyfoolmultimodalllms or with a focus on scientific disciplines such as mathematics wang2024measuringmultimodalmathematicalreasoning ; lu2024mathvistaevaluatingmathematicalreasoning science li2023scigraphqalargescalesyntheticmultiturn ; liang2024scemqascientificcollegeentrance , and programming song2025csbench .
+Specifically, most multimodal benchmarks are designed either from a general perspective Liu et al.
+(2024b; 2023); Meng et al.
+(2024); Han et al.
+(2024); Qian et al.
+(2024) or with a focus on scientific disciplines such as mathematics Wang et al.
+(2024a); Lu et al.
+(2024) science Li and Tajbakhsh (2023); Liang et al.
+(2024), and programming Song et al.
+(2025).
 Unlike STEM fields that employ "vertical reasoning": a focused sequential process using logical deduction and experimental analysis to arrive at singular correct answers, the Humanities and Social Sciences (HSS) emphasize "horizontal reasoning," requiring connections across different contexts and generating multiple valid interpretations rather than single solutions.
@@
 Our contributions can be summarized as follows:
-Figure 2: An example of cross-modal knowledge transfer issues in MLLMs within the HSS domain.
-They struggle to associate "Business Penmanship" font knowledge with relevant images or recognize fonts in image text.
   • First, we introduced HSSBench, a novel dataset specifically designed for the HSS domain, which encompasses 6 distinct categories and 45 major types of HSS tasks.
@@
 This work establishes a foundation for future MLLM research focusing on HSS and serves as a benchmark for future studies in this field.
+Figure 3: Our pipeline to build HSSBench.
 2 HSSBench
@@
 Our dataset focuses on six high-interest categories within the HSS domain: geography, art, culture, social sciences, history, and economics.
-Textual data are the primary medium for disseminating knowledge in the HSS domain, with almost no data in pictorial form.
+Textual data remain the primary medium for disseminating knowledge in the HSS domain, while pictorial data, though less abundant and more challenging to collect, provide valuable complementary information.
 This presents significant challenges for experts in the construction of datasets.
@@
 Data that do not conform to the construction logic are removed.
-2.1 Dataset Preparation, Stage \@slowromancapi@
+2.1 Dataset Preparation, Stage I
 At this stage, we focus on collecting data on target disciplines, covering texts and illustrations.
 Through systematic organization and strict screening, we lay a solid data foundation for the subsequent development of high-quality questions and answers.
-Figure 3: Our pipeline to build HSSBench.
-Data Selection and Collection with Experts
+Data Selection and Collection with Experts.
 During the data preparation for HSSBench, ensuring data diversity is crucial. For image and text selection, we adopt a combined approach involving experts and a multi-agent framework.
@@
 Each expert reviews materials in their field, extracting relevant text passages and images, eliminating redundant information and standardizing the data format.
-Networked Information-Aggregation Agent
+Networked Information-Aggregation Agent.
 Inspired by the expert-led data collection process, we designed the Networked Information-Aggregation Agent to mimic their workflow.
@@
 Finally, domain experts review the curated data to ensure professionalism and reliability.
-2.2 Dataset Construction, Stage \@slowromancapii@
+2.2 Dataset Construction, Stage II
 At this stage, we developed a multidisciplinary visual-question generation pipeline comprising two key phases: Experts Construction and Multi-Agent Construction.
@@
 Dataset Construction with experts.
-We require each question to be presented in a multiple-choice format, with one correct answer and three distractors that were plausible but ultimately incorrect.
 At this stage, experts have two main responsibilities: revising existing questions and creating entirely new ones.
@@
 This Multi-Agent approach enables large-scale generation of high-quality question datasets efficiently and with minimal human intervention.
-2.3 Validation, Stage \@slowromancapiii@
+2.3 Validation, Stage III
 This subsection outlines the stringent validation procedures implemented to ensure the quality and relevance of the data used in model evaluation.
@@
 To meet these two requirements, the Validation Agent assesses the extent to which the question depends on the image.
-If the image is deemed unnecessary, the question is sent back to the Stage \@slowromancapii@ for revision; if it still fails to meet the requirements after multiple iterations, the question is discarded.
+If the image is deemed unnecessary, the question is sent back to the Stage II for revision; if it still fails to meet the requirements after multiple iterations, the question is discarded.
 Expert Validation.
@@
 In our VGP, we leverage GPT-4o and GPT-4.1 due to their outstanding capabilities.
-Throughout all stages of LLM, we employed the Chain-of-Thought prompting strategy.
+Throughout all stages of LLM, we employed the Chain-of-Thought(COT) prompting strategy.
 For more detailed information on VGP, please refer to Appendix A.
 2.5 Data Statistics
-Our dataset comprises a total of 13,152 questions spanning six categories and 45 types.
+HSSBench consists of 13,152 multiple-choice questions distributed across six major categories—Economy, Art, Culture, Social Sciences, History, and Geography—and further divided into 45 specialized subtypes.
+We require each question to be presented in a multiple-choice format with one correct answer and several distractors that are plausible but ultimately incorrect.
+Although some original questions in the HSS domain naturally have multiple correct answers, these multi-answer questions have been reformulated into single-answer questions for consistency.
+This was achieved by combining the multiple correct options into a single choice.
+This approach preserves the original content and complexity of the questions while conforming to a uniform single-answer multiple-choice format.
 It supports evaluation in six different languages.
+Our initial data came from multiple countries and languages, with questions originally created by domain experts in their source languages (examples in Appendix D).
+We then used LLM-based translation models to translate the questions into five other languages.
+All translations were carefully reviewed and validated by bilingual experts to ensure linguistic accuracy and cultural appropriateness, maintaining semantic consistency while respecting cultural nuances.
 The statistical data of HSSBench is illustrated in Figure 4, with a detailed description of the dataset provided in Appendix B.
+Appendix B.1 reports human expert accuracy across all categories.
+Appendix B.2 provides an in-depth taxonomy of categories and subtypes.
+Appendix B.3 also includes an analysis of the most frequent content words per category.
+Appendix B.4 details the composition of the dataset in terms of contributions of human experts versus automated agents, together with quality validation through model accuracy comparisons.
 3 Experiment
@@
 Models.
-For open-source models, we selected Qwen2.5-VL-3B/7B/32B/72B Qwen2.5-VL , Qwen2-VL-72B, QVQ-72B-Preview Qwen2-VL , Deepseek-VL2-Tiny wu2024deepseekvl2mixtureofexpertsvisionlanguagemodels , MiniCPM-o-2.6 yao2024minicpm , mPLUG-Owl3-2B/7B ye2024mplugowl3longimagesequenceunderstanding , Llava-onevision-7B li2024llavaonevisioneasyvisualtask , Llama3-llava-next-8b li2024llavanext-strong , InternVL3-8B zhu2025internvl3exploringadvancedtraining , InternVL2.5-8B-MPO chen2024expanding , Phi-3.5-Vision-Instruct abdin2024phi3technicalreporthighly , Janus-Pro chen2025janusprounifiedmultimodalunderstanding , Llava1.5 liu2024improvedbaselinesvisualinstruction .
-For closed-source commercial models, we utilized GPT-4o/4.1/4.1-mini/4.1-nano openai2024gpt4ocard .
+For open-source models, we selected Qwen2.5-VL-3B/7B/32B/72B Bai et al.
+(2025), Qwen2-VL-72B Wang et al.
+(2024c), QVQ-72B-Preview Wang et al.
+(2024b), Deepseek-VL2-Tiny Wu et al.
+(2024c), MiniCPM-o-2.6 Yao et al.
+(2024), mPLUG-Owl3-2B/7B Ye et al.
+(2024), Llava-onevision-7B Li et al.
+(2024b), Llama3-llava-next-8b Li et al.
+(2024a), InternVL3-8B Zhu et al.
+(2025), InternVL2.5-8B-MPO Chen et al.
+(2024b), Phi-3.5-Vision-Instruct Abdin et al.
+(2024), Janus-Pro Chen et al.
+(2025), Llava1.5 Liu et al.
+(2024a).
+For closed-source commercial models, we utilized GPT-4o/4.1/4.1-mini/4.1-nano OpenAI et al.
+(2024).
 Evaluation Settings.
 1) We conducted evaluations using two types of prompts to examine the MLLMs performance under different prompting strategies.
-One approach involved prompting the model to directly output the answer without any intermediate reasoning, while the other required the model to generate a chain of thought before providing the final answer.
+One approach involved prompting the model to directly output the answer without any intermediate reasoning, while the other required the model to generate a COT before providing the final answer.
 2) Additionally, we tested each question in six different language versions to investigate how the MLLMs’ performance varies across languages for the same question.
@@
 This allowed us to assess whether the model can correctly answer questions without any explicit hints.
-4) For the HSSBench evaluation, we employed two assessment methods and selected GPT-4o as a representative closed-source model and Qwen2.5-7B as a representative open-source model.A detailed comparative analysis of both models is provided in Appendix C.5.
+Detailed prompt settings are provided in Appendix C.8.
+4) For the HSSBench evaluation, we employed two assessment methods and selected GPT-4o as a representative closed-source model and Qwen2.5-7B as a representative open-source model.
+A detailed comparative analysis of both models is provided in Appendix C.9.
 3.2 Results
@@
 and Dt.
-represent two types of prompts: Chain-of-Thought (COT) prompts and direct response prompts.
+represent two types of prompts: COT prompts and direct response prompts.
 C.
@@
 In contrast, some open-source models show considerably lower performance.
-Although data in the HSS domain was generally considered common knowledge, our study demonstrates that from a multimodal perspective, these models require more diverse training data and targeted fine-tuning to effectively address related tasks.
 We also observe that model performance varies by language.
@@
 Human 94.14 - 93.06 - 92.99 - 94.44 - 93.84 - 95.53 - 93.83 - 93.83 -
-Open-source LLM (Scale <<< 10B)
+Open-source LLM (Scale << 10B)
 Qwen2.5-VL-3B-Instruct 35.25 11.74 28.56 16.36 34.61 1.85 39.46 11.50 36.23 11.07 34.55 3.62 29.01 9.94 34.99 9.33
@@
 Llava1.5 12.75 4.03 7.38 2.68 10.74 3.36 7.38 5.37 10.74 4.03 10.74 6.04 8.06 4.03 9.96 4.25
-Open-source LLM (Scale >>> 10B)
+Open-source LLM (Scale >> 10B)
 Qwen2.5-VL-32B-Instruct 52.48 21.33 52.79 6.67 38.94 8.00 53.87 24.00 57.03 26.67 39.20 3.33 48.38 15.89 50.75 15.00
@@
 GPT-4.1-nano 33.21 30.74 39.71 41.44 38.70 7.83 37.52 28.14 34.01 30.74 35.83 20.36 36.33 21.12 35.83 26.22
-Table 1: Scores (%) of MLLMs on HSSBench (EN-\@slowromancapi@). The highest and second highest scores are marked in blue and green, respectively.
+Table 1: Scores (%) of MLLMs on HSSBench (EN-I). The highest and second highest scores are marked in blue and green, respectively.
 Model performance on different categories.
-Among the six categories evaluated on HSSBench (EN-\@slowromancapi@), the economic-related tasks consistently emerge as the most challenging for the models.
+Among the six categories evaluated on HSSBench (EN-I), the economic-related tasks consistently emerge as the most challenging for the models.
 The average score of all models in this category is the lowest, indicating that addressing economic problems requires a deep understanding of various economic theories and the ability to apply them in complex reasoning.
@@
 Comparison between Direct Answer and COT Prompting.
-Table 1 demonstrates that the performance of various MLLMs on HSSBench using direct answer prompts and COT prompts exhibits subtle differences.
-COT prompts do not always yield positive effects, and certain models perform better with direct answer prompts.
-This suggests that extending the reasoning process might mislead the models.
-In HSS tasks, the COT method exhibits limitations for certain models.
-It not only fails to deepen the models’ understanding but may also cause them to focus on irrelevant information, thereby interfering with attention to critical knowledge.
+Table 1 shows subtle differences in MLLMs’ performance on HSSBench when using direct answer versus COT prompting.
+Notably, COT does not always help; some models perform better with direct answers, indicating that longer reasoning can mislead them.
 Specifically, COT prompts exacerbate hallucination issues in certain models, where reasoning flaws in textual analysis and misinterpretation of visual inputs lead to the generation of incorrect background knowledge during step-by-step analysis.
@@
 Comparison between Multiple-choices and Open-ended.
-We selected questions from the dataset that can be reformulated as open-ended to evaluate model performance without the guidance of provided answer options.
-The results indicate that HSS tasks pose significant challenges for current MLLMs, with only a few models achieving accuracy above 15% in the open-ended setting.
-This result aligns with our expectations: even experts find it difficult to provide the correct answer without predefined options and background knowledge.
-The presence of answer options offers models valuable prior knowledge, effectively constraining the scope of possible answers.
-In contrast, open-ended questions remove this constraint, causing the model’s reasoning to become highly divergent and often leading to answers that deviate substantially from the ground truth.
-This suggests that models primarily rely on shallow visual attributes, such as object size, type, or motion, while overlooking deeper symbolic information embedded in cultural context, academic theories, or spatiotemporal cues.
-These observations highlight two key limitations of current MLLMs: first, a lack of comprehensive understanding and memory of real-world knowledge; second, difficulty in effectively integrating visual features with textual information.
-Moreover, the performance degradation observed when introducing chain-of-thought prompting further underscores the models’ struggles with internalizing and retrieving complex knowledge.
+We reformulated some questions as open-ended to test models without answer options.
+Results reveal that HSS tasks remain very challenging: only a few models exceed 15% accuracy.
+This matches expectations, as even experts find these questions difficult without options or background cues.
+Answer choices provide prior knowledge that narrows the answer space.
+Without them, models’ reasoning becomes highly divergent, often drifting far from correct answers.
+This suggests models mainly rely on shallow visual features (e.g., size, type, motion) but miss deeper symbolic information like cultural context or spatiotemporal cues.
+These findings highlight two key limitations: insufficient real-world knowledge and weak integration of visual and textual modalities.
+The performance drop with COT further reflects models’ struggles to internalize and retrieve complex knowledge.
+Improving model capabilities on HSS tasks remains a critical challenge.
 The role of accurate answers.
-In multiple-choice settings, the presence of answer options inevitably provides additional cues that can influence model performance.
-To further investigate whether models truly understand the questions rather than relying on superficial cues, we introduced confusing options, specifically, the choice "None of the above," while keeping the correct answer unchanged.
-This experiment was conducted on a subset of 150 samples randomly selected from each category.
-The results are illustrated in Table 5.
-For lower-performing models, the introduction of confusing options led to a noticeable decline in performance.
-Upon closer examination of the models’ reasoning processes after adding these distractors, two main patterns emerged.
-In some cases, models were directly swayed by the confusing option and selected "None of the above" incorrectly.
-More intriguingly, in other cases, although the models did not explicitly choose the confusing option, their reasoning became less coherent.
-This behavior indicates that current models struggle to accurately assess the relative credibility of different answer choices and to make reliable judgments when faced with conflicting information.
-Consequently, their performance deteriorates in the presence of carefully designed distractors.
+To test if models truly understand questions rather than exploiting options, we added confusing choices like “None of the above” while keeping the correct answer unchanged, sampling 150 questions per category (results in Table 11, Appendix C.3).
+Lower-performing models’ accuracy dropped noticeably.
+Analysis revealed two patterns: some models were misled to select the confusing option; others, though not choosing it, showed incoherent reasoning, failing to filter out wrong choices and sometimes excluding the correct answer altogether.
+This indicates models struggle to assess the relative credibility of options and make reliable judgments, exposing their vulnerability to distractors and raising concerns about robustness.
 Visual Information Extraction.
@@
 In the first experiment, we used GPT-4.1 to generate detailed textual descriptions of the images.
-In the second experiment, we invited domain experts to produce comprehensive and precise annotations for each image.
-The experimental results are summarized in Table 6.
-Specifically, we converted the content of images into textual descriptions and provided only these texts, along with the questions and answer options, to the models.
-The accuracy of most models declined compared to when the full images were available, indicating that direct visual input contains critical information that is partially lost in the text conversion process.
-However, considering that even when models receive the full images, they may not fully comprehend certain fine-grained visual details, we further involved domain experts to produce detailed annotations of the images.
-These expert annotations distilled key visual information into precise textual descriptions.
-When provided with these expert-annotated texts instead of the images, the models’ accuracy notably improved relative to the plain text conversion scenario, despite the absence of the original images.
-Interestingly, some models even surpassed their previous upper-bound performance in domains such as Culture and Social Sciences.
-This improvement suggests that expert annotations effectively guide the models to focus on critical visual cues by translating them into semantically rich and relevant textual information, thereby facilitating stronger connections with domain knowledge.
-These findings corroborate our earlier observations on open-ended questions, highlighting inherent limitations in current MLLMs’ abilities to retrieve and interpret visual information comprehensively.
-4 Limitations
-Our HSSBench is designed to evaluate MLLMs’ capabilities through HSS reasoning tasks.
-The type within our dataset reflects our effort to collect data from as many perspectives as possible.
-However, we acknowledge that the scope of knowledge covered is inherently limited by the contributors’ expertise and cannot encompass all aspects of HSS.
-Moreover, HSS currently lacks a universally accepted definition.
-Our evaluation is based on six major categories that, while comprehensive, do not cover the full spectrum of HSS-related fields.
-In addition, some data points involve recent social events and the latest economic knowledge.
-We recognize that such data are time sensitive.
-Although the answers are accurate at present, future developments may lead to different interpretations.
-For these items, we impose temporal and spatial constraints to ensure their correctness within the current context.
-Although these limitations are relevant, we believe that they extend beyond the scope and original intent of our dataset. We hope that future work will address these challenges in more detail.
+In the second experiment, we invited domain experts to produce comprehensive and precise annotations for each image (results in Table 12, Appendix C.4).
+Most models’ accuracy dropped when images were replaced by GPT-generated texts, showing that direct visual input contains critical details lost in conversion.
+However, with expert annotations, accuracy improved noticeably despite no image access.
+Some models even surpassed previous performance ceilings in Culture and Social Sciences, as expert texts helped focus on crucial visual cues and better link to domain knowledge.
+These results confirm earlier observations that current MLLMs have inherent limitations in retrieving and understanding visual information fully.
+They underscore the importance of improving models’ abilities to extract and integrate key visual features for HSS tasks.
+Comparative Analysis of HSSBench with Related Benchmarks.
+We compared HSSBench with other benchmarks that include some HSS data, such as CMMMU, MME, and MMMU, as well as with STEM benchmarks (details in Appendix C.5 and C.6).
+Across HSS benchmarks, relative model performance trends are consistent, but HSSBench yields lower accuracy, indicating it is more challenging and better captures HSS complexity.
+Compared to STEM benchmarks, models perform much better on STEM tasks, highlighting the unique difficulties of HSS domains that require nuanced cultural and social understanding.
+These comparisons demonstrate HSSBench’s value in driving progress on underexplored HSS challenges.
+Evaluation of Retrieval-Augmented Generation on HSSBench.
+We tested Retrieval-Augmented Generation (RAG) by integrating external knowledge from Wikipedia and HSS documents with several smaller MLLMs under direct and COT prompting (results in Tables 15 and 16, Appendix C.7).
+Contrary to expectations, RAG did not consistently improve accuracy.
+Some modest gains appeared in specific domains or models, but overall performance was often similar or worse than direct prompting alone.
+This suggests that general retrieval corpora and simple integration methods are insufficient for the nuanced knowledge HSS tasks demand.
+The limited effectiveness of RAG in supporting MLLMs for complex multi-hop HSS tasks stems from misalignment between general-purpose corpora and HSS knowledge, as well as deficiencies in current retrieval and integration methods.
+Moreover, MLLMs struggle to internalize and transfer newly retrieved domain knowledge.
+These challenges highlight the need for specialized HSS retrieval resources, RAG techniques tailored to HSS, and enhanced mechanisms for MLLMs to acquire, transfer, and apply HSS-specific knowledge.
 4 Related Work
 In recent years, significant progress has been made in the development of multimodal benchmarks and methodologies.
-Numerous datasets have emerged that assess models from various perspectives, which can be broadly categorized into Generation Benchmarks liu2024mmbench ; liu2023visualinstructiontuning ; meng2024vgavisionguiassistant ; han2024instinctivebiasspuriousimages ; qian2024easyfoolmultimodalllms ; xu2023lvlmehubcomprehensiveevaluationbenchmark ; yin2023lammlanguageassistedmultimodalinstructiontuning ; zeng2023matterstraininggpt4stylelanguage ; wu2024qbenchbenchmarkgeneralpurposefoundation ; luo-etal-2024-codis , Reasoning Benchmarks wang2024measuringmultimodalmathematicalreasoning ; lu2024mathvistaevaluatingmathematicalreasoning ; li2023scigraphqalargescalesyntheticmultiturn ; liang2024scemqascientificcollegeentrance ; song2025csbench , and Application Benchmarks fan2022minedojobuildingopenendedembodied ; rawles2023androidwildlargescaledataset ; chen2023endtoendembodieddecisionmaking ; sermanet2023robovqamultimodallonghorizonreasoning ; chen2024egoplanbenchbenchmarkingmultimodallarge ; koh2024visualwebarenaevaluatingmultimodalagents ; you2024ferretuigroundedmobileui .
-Generation benchmarks cover a wide range of tasks, with the core objective of comprehensively evaluating the multi-dimensional performance of MLLMs.
-They not only include reasoning challenges but also address areas related to humanities and social sciences.
+Numerous datasets have emerged that assess models from various perspectives, which can be broadly categorized into Generation Benchmarks Liu et al.
+(2024b; 2023); Meng et al.
+(2024); Han et al.
+(2024); Qian et al.
+(2024); Xu et al.
+(2023); Yin et al.
+(2023); Zeng et al.
+(2023); Wu et al.
+(2024a); Luo et al.
+(2024); Feng et al.
+(2025); Wang et al.
+(2025e), Reasoning Benchmarks Wang et al.
+(2024a); Lu et al.
+(2024); Li and Tajbakhsh (2023); Liang et al.
+(2024); Song et al.
+(2025), and Application Benchmarks Fan et al.
+(2022); Rawles et al.
+(2023); Chen et al.
+(2023); Sermanet et al.
+(2023); Chen et al.
+(2024a); Koh et al.
+(2024); You et al.
+(2024); Gao et al.
+(2025); Kang et al.
+(2026); Zheng et al.
+(2026); He et al.
+(2026).
+Generation benchmarks cover a wide range of tasks, often including not only reasoning challenges but also content related to the humanities and social sciences.
+Their main objective is to provide a comprehensive assessment of the performance of MLLMs in various dimensions.
 In contrast, reasoning benchmarks are designed primarily around mathematical and scientific problems, seeking to rigorously assess the capabilities of MLLM in logical and analytical reasoning.
-Among the various evaluation formats, multiple-choice datasets are the most prevalent, owing to their simplicity in evaluation and ease of comparison zeng2023matterstraininggpt4stylelanguage .
-The proliferation of these data sets has significantly accelerated research progress, serving both as training resources and as tools for assessing the multifaceted competencies of MLLMs.
-Advances in data set construction are particularly exciting, ranging from human-annotated high-quality datasets romero2024cvqaculturallydiversemultilingualvisual to those generated through pipelines built in LLM chandrasegaran2024hourvideo1hourvideolanguageunderstanding .
+Among the various evaluation formats, multiple-choice datasets are the most prevalent, owing to their simplicity in evaluation and ease of comparison Zeng et al.
+(2023).
+The proliferation of these data sets has significantly accelerated research progress, serving both as training resources and as tools for assessing the multifaceted competencies of MLLMs Wang et al.
+(2025f); Liu et al.
+(2025b); Luo et al.
+(2026); Guan et al.
+(2026); Fu et al.
+(2026); Wu et al.
+(2025; 2024b); Feng et al.
+(2026); Li et al.
+(2026; 2025c; 2025d); Hu et al.
+; Zhang et al.
+(2025c; b; a); Ding et al.
+(2025); Li et al.
+(2025b); Wang et al.
+(2026); Ma et al.
+(2026); Liu et al.
+(2025a); Bi et al.
+(2026); Wang et al.
+(2025d; c; b; a); Liang et al.
+(2025); Ma et al.
+(2025a; b); Li et al.
+(2025a); Yu et al.
+(2025); Feng et al.
+(2024; 2023).
+Advances in data set construction are particularly exciting, ranging from human-annotated high-quality datasets Romero et al.
+(2024) to those generated through pipelines built in LLM Chandrasegaran et al.
+(2024).
 Despite these achievements, several critical limitations remain.
@@
 They can select entries to view detailed previews of the entries.
+A.2 Information about Experts
+Name Languages Professional Background
+Expert 1 Chinese, English Computer Science, Art, Economy
+Expert 2 Chinese, Japanese, English Computer Science, Culture
+Expert 3 English, Chinese Computer Science, Geography
+Expert 4 Chinese, English Economy, History
+Expert 5 Chinese, English, Japanese Language and Linguistics, History, Culture
+Expert 6 Chinese, English, Russian Language and Linguistics, Social Science
+Expert 7 Chinese, English, Arabic Language and Linguistics, Social Science
+Expert 8 Chinese, English, Arabic Economy, Social Science
+Expert 9 Chinese, English Geography
+Expert 10 Chinese, English, French Economy, Art
+Expert 11 English, Chinese, Japanese Art, Culture
+Expert 12 Chinese, English Computer Science, Geography
+Expert 13 Chinese, English Geography, History
+Expert 14 Chinese, English, Spanish Computer Science, Geography
+Table 2: Information about the experts involved in dataset construction.
+The team of experts who contributed to the creation of HSSBench comes from a variety of cultural and linguistic backgrounds, as well as interdisciplinary academic fields spanning the humanities and social sciences.
+This diversity is reflected in their linguistic fluency, professional expertise, and international academic experiences.
+Table 2 summarizes the profiles of the experts involved in the dataset construction.
+Many of the experts have international academic experiences and interdisciplinary training across various humanities and social science domains.
+While individual cultural perspectives may naturally influence the emphasis or framing of certain questions, we consider this diversity a strength rather than a limitation.
+It enriches the dataset by incorporating multiple viewpoints and insights, which is essential for a benchmark designed to be multilingual and cross-cultural.
+To mitigate potential cultural bias, we carefully curated the source materials provided to the experts during the initial stage of dataset construction.
+These reference materials originate from multiple countries and languages, ensuring that the constructed questions and answers reflect widely accepted, global knowledge rather than culturally subjective viewpoints.
+Furthermore, during the data validation and filtering stages, we rigorously excluded content that could be culturally biased or inconsistent with universal values.
+In summary, our approach balances the preservation of valuable cultural diversity with the need to maintain fairness and universality in the dataset.
+This careful design ensures that HSSBench serves as a robust and inclusive benchmark for evaluating multilingual large language models (MLLMs) across different languages and cultural contexts.
 Appendix B Dataset Details
@@
 Figure 6: Word Cloud in HSSBench per category.
+B.4 Composition and Quality Comparison Between Human and Automated Agents
+We provide a detailed breakdown of the data sources across the six domains in HSSBench, as shown in Table 5.
+Category Human Expert Automated Agent Total Samples
+Art 984 1,204 2,188
+Geography 2,671 891 3,562
+History 1,946 550 2,496
+Economy 974 459 1,433
+Social Science 487 947 1,434
+Culture 652 1,387 2,039
+Table 5: Distribution of samples constructed by human experts and automated agents across different domains in HSSBench.
+To further validate the quality of data generated by both human experts and automated agents, we evaluated the performance of GPT4.1-mini on the respective subsets.
+The accuracy results are summarized in Table 6.
+Category Human Expert Acc. Automated Agent Acc.
+Art 42.32 45.43
+Geography 48.82 46.52
+History 45.59 49.12
+Economy 58.87 57.67
+Social Science 43.15 46.95
+Culture 49.11 47.41
+Table 6: Model accuracy (%) of GPT4.1-mini on human-expert-generated and automated-agent-generated subsets across different domains.
+The comparable accuracy results indicate that data generated by automated agents, after undergoing multi-round expert verification, maintain a quality level consistent with that of human expert contributions.
+These clarifications and statistics have been incorporated into the revised manuscript to enhance transparency regarding dataset construction and quality assurance.
 Appendix C Experimental details
@@
 Human 94.14 - 93.06 - 92.99 - 94.44 - 93.84 - 95.53 - 93.83 - 93.83 -
-Open-source LLM (Scale <<< 10B)
+Open-source LLM (Scale << 10B)
 Qwen2.5-VL-3B 29.84 11.27 29.44 17.76 27.83 5.09 35.22 10.00 26.91 8.61 26.78 7.24 29.01 9.94 34.99 9.33
@@
 Llava1.5 3.36 2.68 10.07 5.37 3.36 2.01 8.05 4.70 8.05 1.34 15.44 8.05 8.06 4.03 9.96 4.25
-Open-source LLM (Scale >>> 10B)
+Open-source LLM (Scale >> 10B)
 Qwen2.5-VL-32B 52.54 16.00 39.38 8.00 39.87 10.00 53.09 26.67 57.52 24.67 44.92 10.00 48.38 15.89 50.75 15.00
@@
 GPT-4.1-nano 33.18 20.66 34.96 36.04 45.61 14.75 38.42 15.58 32.69 20.49 36.47 18.55 36.33 21.12 35.83 26.22
-Table 4: Scores (%) of MLLMs on HSSBench (EN-\@slowromancapii@).
+Table 7: Scores (%) of MLLMs on HSSBench (EN-II).
 (a) Effect under multiple-choice question setting
@@
 Figure 7: Partial MLLMs’ results under CoT prompt settings.
-C.2 Details of Contrast Experiment
-Table 5 and Figure 8 provides a detailed presentation of the experimental results after adding a confounding option.
-We sampled 900 data points.
-"Conf." indicates that, in addition to the given options, an extra option—"None of the above answers is correct"—was added.
-The model’s output performance was then compared under these conditions.
+C.2 Results in different languages
+Table 1 and Table 7 present the test results on English-language data.
+To analyze the impact of different languages on the model, we also conducted large-scale experiments using Chinese-language data.
+The final experimental results are shown in Table 8 and Table 9 below.
+Model Geography Economics Culture Social Sciences History Art All
+Ct.C. Ct.O. Ct.C. Ct.O. Ct.C. Ct.O. Ct.C. Ct.O. Ct.C. Ct.O. Ct.C. Ct.O. Dr.C. Dr.O. Ct.C. Ct.O.
+Random 24.93 0.00 21.92 0.00 25.00 0.00 24.90 0.00 24.91 0.00 25.00 0.00 24.62 0.00 24.62 0.00
+Human 94.14 - 93.06 - 92.99 - 94.44 - 93.84 - 95.53 - 93.83 - 93.83 -
+Open-source LLM (Scale << 10B)
+Qwen2.5-VL-3B 40.49 10.80 28.89 19.81 36.30 0.93 42.43 9.50 43.91 11.89 33.85 4.98 35.45 11.79 35.45 9.65
+Qwen2.5-VL-7B 48.43 15.96 35.00 25.00 27.60 2.78 49.36 19.50 50.54 15.98 32.90 6.79 43.12 13.17 41.86 14.24
+Llava-onevision-7b 39.68 4.69 30.80 10.38 32.96 1.85 39.46 7.00 32.97 4.51 35.03 2.71 37.89 4.98 35.62 5.13
+Llama3-llava-next-8b 27.29 3.29 19.82 3.77 33.55 0.93 32.89 2.50 26.84 3.28 29.48 2.26 31.18 5.21 28.35 2.68
+InternVL3-8B 47.45 11.74 35.37 17.92 38.21 9.26 53.47 14.00 50.97 18.44 39.94 10.41 44.81 14.70 44.92 13.71
+InternVL2.5-8B-MPO 41.14 11.74 36.11 23.58 36.20 6.94 48.09 16.50 47.63 6.94 38.35 10.86 42.72 14.01 41.45 14.55
+Phi-3.5-vision-instruct 22.29 6.10 25.64 4.25 30.16 2.31 23.55 1.00 23.35 4.92 28.30 1.81 34.08 6.74 25.35 3.45
+Janus-Pro 28.10 11.27 21.67 8.49 44.16 6.48 34.79 8.50 29.13 14.34 40.49 14.48 33.50 11.94 32.81 10.72
+mPLUG-Owl3-2B 27.20 1.88 29.43 0.90 33.74 2.30 29.78 3.00 26.76 2.87 29.43 4.52 32.25 1.90 28.25 2.58
+mPLUG-Owl3-7B 35.01 5.63 34.23 3.15 26.39 1.84 39.26 2.50 26.08 2.46 34.23 5.43 34.22 4.71 31.77 3.49
+MiniCPM-o-2.6 26.97 7.69 19.50 4.88 35.21 6.25 31.23 3.90 30.27 6.76 30.20 5.98 1.29 6.83 29.04 5.94
+Open-source LLM (Scale >> 10B)
+Qwen2.5-VL-32B 58.45 19.33 49.86 7.33 40.90 5.33 56.49 22.67 64.33 24.67 41.62 14.67 51.60 14.22 52.86 15.67
+Qwen2.5-VL-72B 61.65 12.67 52.02 33.33 43.60 10.67 61.04 14.19 68.08 18.12 43.76 16.00 60.04 21.96 55.94 17.51
+QVQ-72B-Preview 19.82 8.00 21.10 18.00 30.80 3.33 25.41 6.76 27.85 10.74 25.90 6.67 22.95 9.14 24.82 8.92
+Qwen2-VL-72B 21.28 19.30 38.25 24.07 21.25 5.66 27.79 12.96 25.69 32.08 27.65 18.87 59.11 17.79 25.59 18.82
+Closed-source LLM
+GPT-4.1mini 47.53 35.21 55.55 55.41 44.43 18.43 44.07 34.17 47.92 43.03 42.50 30.32 44.19 24.54 46.78 36.32
+Table 8: Scores (%) of MLLMs on HSSBench (ZH-I).
+Model Geography Economics Culture Social Sciences History Art All
+Dr.C.g Dr.O.g Dr.C.g Dr.O.g Dr.C.g Dr.O. Dr.C.g Dr.O.g Dr.C.g Dr.O.g Dr.C.g Dr.O.g Dr.C.g Dr.O.g Ct.C.g Ct.O.g
+Random 24.93 0.00 21.92 0.00 25.00 0.00 24.90 0.00 24.91 0.00 25.00 0.00 24.62 0.00 24.62 0.00
+Human 94.14 - 93.06 - 92.99 - 94.44 - 93.84 - 95.53 - 93.83 - 93.83 -
+Open-source LLM (Scale << 10B)
+Qwen2.5-VL-3B 39.51 12.21 29.26 13.21 30.94 7.87 40.03 14.50 38.36 11.48 30.03 11.76 35.45 11.79 35.45 9.65
+Qwen2.5-VL-7B 52.52 17.84 33.97 20.75 27.60 4.63 51.13 15.00 52.13 11.48 31.17 9.95 43.12 13.17 41.86 14.24
+Llava-onevision-7b 42.20 6.57 30.21 7.55 36.59 1.39 42.93 5.00 35.61 4.10 36.12 5.43 37.89 4.98 35.62 5.13
+Llama3-llava-next-8b 29.61 4.69 19.97 6.60 38.80 5.09 34.37 5.00 28.86 3.28 34.54 6.33 31.18 5.21 28.35 2.68
+InternVL3-8B 46.44 18.78 34.27 20.75 40.08 6.94 52.90 18.00 51.05 14.34 40.19 9.95 44.81 14.70 44.92 13.71
+InternVL2.5-8B-MPO 42.26 18.78 31.32 14.15 40.72 7.87 48.87 17.50 49.11 12.30 40.78 14.03 42.72 14.01 41.45 14.55
+Phi-3.5-vision-instruct 29.92 6.57 27.49 7.08 45.58 7.87 32.89 4.50 29.71 6.56 40.68 7.69 34.08 6.74 25.35 3.45
+Janus-Pro 27.29 9.86 18.50 12.74 50.83 14.35 32.18 8.50 29.44 10.25 43.21 15.84 33.50 11.94 32.81 10.72
+mPLUG-Owl3-2B 31.25 0.94 23.73 2.71 35.95 0.46 37.17 3.00 31.89 2.46 33.18 2.71 32.25 1.90 28.25 2.58
+mPLUG-Owl3-7B 35.85 7.51 35.65 6.79 35.70 1.38 40.59 3.00 28.17 4.10 35.65 6.79 34.22 4.71 31.77 3.49
+MiniCPM-o-2.6 0.93 7.55 0.77 8.06 1.42 1.61 1.99 5.41 0.60 20.00 2.42 2.27 1.29 6.83 29.04 5.94
+Open-source LLM (Scale >> 10B)
+Qwen2.5-VL-32B 56.29 20.67 42.83 6.00 41.83 4.00 57.42 18.00 63.00 25.33 42.21 11.33 51.60 14.22 52.86 15.67
+Qwen2.5-VL-72B 68.15 22.67 48.40 30.00 50.07 12.00 63.73 20.27 71.02 31.54 49.02 15.33 60.04 21.96 55.94 17.51
+QVQ-72B-Preview 19.40 6.67 17.90 20.00 29.72 4.00 24.34 11.49 28.13 8.05 23.44 4.67 22.95 9.14 24.82 8.92
+Qwen2-VL-72B 64.42 28.07 41.83 16.67 55.98 7.55 62.00 14.81 68.77 22.64 51.49 16.98 59.11 17.79 25.59 18.82
+Closed-source LLM
+GPT-4.1mini 45.93 26.76 53.31 43.69 46.25 14.29 39.61 20.10 39.26 24.18 42.09 17.65 44.19 24.54 46.78 36.32
+Table 9: Scores (%) of MLLMs on HSSBench (ZH-II).
+Figure 8 illustrates the performance of various models under four different prompt configurations.
+Figure 8: Comparison of Model Performances Across Four Prompt Settings.
+Table 10 and Figure 9 shows the performance of the model on datasets in different languages.
+Our dataset includes data organized in six languages.
+The table below presents the experimental results on a stratified sample of 900 instances.
+C.3 Multiple-Choice Confounding Option Experiment Details
+Model Arabic Chinese English French Russian Spanish
+InternVL3-8B 35.20 49.37 42.38 39.37 38.70 39.21
+Qwen2.5-VL-7B-Instruct 33.74 40.21 38.37 34.08 35.55 34.83
+Qwen2.5-VL-32B-Instruct 46.94 54.79 50.24 47.57 48.51 50.24
+Qwen2.5-VL-72B-Instruct 49.72 55.41 51.28 51.28 48.27 48.83
+QVQ-72B-Preview 33.98 36.89 39.80 37.86 40.77 38.83
+GPT-4.1-mini 41.33 44.89 41.67 41.78 41.44 46.89
+Average 40.82 46.93 43.96 41.99 42.87 43.47
+Table 10: Contrast Scores (%) of MLLMs on HSSBench for six UN languages and six models.
+Figure 9: Contrast Scores (%) of MLLMs on HSSBench for six UN languages and six models.
 Model Geography Art Culture Social Sciences History Economy Total
@@
 Table 11: Contrast Scores (%) with confounding option of MLLMs on HSSBench.
+Table 11 and Figure 10 provides a detailed presentation of the experimental results after adding a confounding option.
+We sampled 900 data points.
+"Conf." indicates that, in addition to the given options, an extra option—"None of the above answers is correct"—was added.
+The model’s output performance was then compared under these conditions.
 (a) Art
@@
 Figure 10: Contrast Scores (%) with confounding option of MLLMs on HSSBench.
+C.4 Visual Information Extraction Experiment Details
 Table 12 and Figure 11 presents the extraction of image information into text (where "De." indicates a detailed description of the image generated directly by GPT-4.1, and "De-H." indicates a detailed explanation provided by a domain expert for each image).
@@
 Figure 11: Contrast Scores (%) about Visual Information Extraction of MLLMs on HSSBench.
-C.3 Results in different languages
-Table 1 and Table 4 present the test results on English-language data.
-To analyze the impact of different languages on the model, we also conducted large-scale experiments using Chinese-language data.
-The final experimental results are shown in Table 7 and Table 8 below.
-Model Geography Economics Culture Social Sciences History Art All
-Ct.C. Ct.O. Ct.C. Ct.O. Ct.C. Ct.O. Ct.C. Ct.O. Ct.C. Ct.O. Ct.C. Ct.O. Dr.C. Dr.O. Ct.C. Ct.O.
-Random 24.93 0.00 21.92 0.00 25.00 0.00 24.90 0.00 24.91 0.00 25.00 0.00 24.62 0.00 24.62 0.00
-Human 94.14 - 93.06 - 92.99 - 94.44 - 93.84 - 95.53 - 93.83 - 93.83 -
-Open-source LLM (Scale <<< 10B)
-Qwen2.5-VL-3B 40.49 10.80 28.89 19.81 36.30 0.93 42.43 9.50 43.91 11.89 33.85 4.98 35.45 11.79 35.45 9.65
-Qwen2.5-VL-7B 48.43 15.96 35.00 25.00 27.60 2.78 49.36 19.50 50.54 15.98 32.90 6.79 43.12 13.17 41.86 14.24
-Llava-onevision-7b 39.68 4.69 30.80 10.38 32.96 1.85 39.46 7.00 32.97 4.51 35.03 2.71 37.89 4.98 35.62 5.13
-Llama3-llava-next-8b 27.29 3.29 19.82 3.77 33.55 0.93 32.89 2.50 26.84 3.28 29.48 2.26 31.18 5.21 28.35 2.68
-InternVL3-8B 47.45 11.74 35.37 17.92 38.21 9.26 53.47 14.00 50.97 18.44 39.94 10.41 44.81 14.70 44.92 13.71
-InternVL2.5-8B-MPO 41.14 11.74 36.11 23.58 36.20 6.94 48.09 16.50 47.63 6.94 38.35 10.86 42.72 14.01 41.45 14.55
-Phi-3.5-vision-instruct 22.29 6.10 25.64 4.25 30.16 2.31 23.55 1.00 23.35 4.92 28.30 1.81 34.08 6.74 25.35 3.45
-Janus-Pro 28.10 11.27 21.67 8.49 44.16 6.48 34.79 8.50 29.13 14.34 40.49 14.48 33.50 11.94 32.81 10.72
-mPLUG-Owl3-2B 27.20 1.88 29.43 0.90 33.74 2.30 29.78 3.00 26.76 2.87 29.43 4.52 32.25 1.90 28.25 2.58
-mPLUG-Owl3-7B 35.01 5.63 34.23 3.15 26.39 1.84 39.26 2.50 26.08 2.46 34.23 5.43 34.22 4.71 31.77 3.49
-MiniCPM-o-2.6 26.97 7.69 19.50 4.88 35.21 6.25 31.23 3.90 30.27 6.76 30.20 5.98 1.29 6.83 29.04 5.94
-Open-source LLM (Scale >>> 10B)
-Qwen2.5-VL-32B 58.45 19.33 49.86 7.33 40.90 5.33 56.49 22.67 64.33 24.67 41.62 14.67 51.60 14.22 52.86 15.67
-Qwen2.5-VL-72B 61.65 12.67 52.02 33.33 43.60 10.67 61.04 14.19 68.08 18.12 43.76 16.00 60.04 21.96 55.94 17.51
-QVQ-72B-Preview 19.82 8.00 21.10 18.00 30.80 3.33 25.41 6.76 27.85 10.74 25.90 6.67 22.95 9.14 24.82 8.92
-Qwen2-VL-72B 21.28 19.30 38.25 24.07 21.25 5.66 27.79 12.96 25.69 32.08 27.65 18.87 59.11 17.79 25.59 18.82
-Closed-source LLM
-GPT-4.1mini 47.53 35.21 55.55 55.41 44.43 18.43 44.07 34.17 47.92 43.03 42.50 30.32 44.19 24.54 46.78 36.32
-Table 7: Scores (%) of MLLMs on HSSBench (ZH-\@slowromancapi@).
-Model Geography Economics Culture Social Sciences History Art All
-Dr.C.g Dr.O.g Dr.C.g Dr.O.g Dr.C.g Dr.O. Dr.C.g Dr.O.g Dr.C.g Dr.O.g Dr.C.g Dr.O.g Dr.C.g Dr.O.g Ct.C.g Ct.O.g
-Random 24.93 0.00 21.92 0.00 25.00 0.00 24.90 0.00 24.91 0.00 25.00 0.00 24.62 0.00 24.62 0.00
-Human 94.14 - 93.06 - 92.99 - 94.44 - 93.84 - 95.53 - 93.83 - 93.83 -
-Open-source LLM (Scale <<< 10B)
-Qwen2.5-VL-3B 39.51 12.21 29.26 13.21 30.94 7.87 40.03 14.50 38.36 11.48 30.03 11.76 35.45 11.79 35.45 9.65
-Qwen2.5-VL-7B 52.52 17.84 33.97 20.75 27.60 4.63 51.13 15.00 52.13 11.48 31.17 9.95 43.12 13.17 41.86 14.24
-Llava-onevision-7b 42.20 6.57 30.21 7.55 36.59 1.39 42.93 5.00 35.61 4.10 36.12 5.43 37.89 4.98 35.62 5.13
-Llama3-llava-next-8b 29.61 4.69 19.97 6.60 38.80 5.09 34.37 5.00 28.86 3.28 34.54 6.33 31.18 5.21 28.35 2.68
-InternVL3-8B 46.44 18.78 34.27 20.75 40.08 6.94 52.90 18.00 51.05 14.34 40.19 9.95 44.81 14.70 44.92 13.71
-InternVL2.5-8B-MPO 42.26 18.78 31.32 14.15 40.72 7.87 48.87 17.50 49.11 12.30 40.78 14.03 42.72 14.01 41.45 14.55
-Phi-3.5-vision-instruct 29.92 6.57 27.49 7.08 45.58 7.87 32.89 4.50 29.71 6.56 40.68 7.69 34.08 6.74 25.35 3.45
-Janus-Pro 27.29 9.86 18.50 12.74 50.83 14.35 32.18 8.50 29.44 10.25 43.21 15.84 33.50 11.94 32.81 10.72
-mPLUG-Owl3-2B 31.25 0.94 23.73 2.71 35.95 0.46 37.17 3.00 31.89 2.46 33.18 2.71 32.25 1.90 28.25 2.58
-mPLUG-Owl3-7B 35.85 7.51 35.65 6.79 35.70 1.38 40.59 3.00 28.17 4.10 35.65 6.79 34.22 4.71 31.77 3.49
-MiniCPM-o-2.6 0.93 7.55 0.77 8.06 1.42 1.61 1.99 5.41 0.60 20.00 2.42 2.27 1.29 6.83 29.04 5.94
-Open-source LLM (Scale >>> 10B)
-Qwen2.5-VL-32B 56.29 20.67 42.83 6.00 41.83 4.00 57.42 18.00 63.00 25.33 42.21 11.33 51.60 14.22 52.86 15.67
-Qwen2.5-VL-72B 68.15 22.67 48.40 30.00 50.07 12.00 63.73 20.27 71.02 31.54 49.02 15.33 60.04 21.96 55.94 17.51
-QVQ-72B-Preview 19.40 6.67 17.90 20.00 29.72 4.00 24.34 11.49 28.13 8.05 23.44 4.67 22.95 9.14 24.82 8.92
-Qwen2-VL-72B 64.42 28.07 41.83 16.67 55.98 7.55 62.00 14.81 68.77 22.64 51.49 16.98 59.11 17.79 25.59 18.82
-Closed-source LLM
-GPT-4.1mini 45.93 26.76 53.31 43.69 46.25 14.29 39.61 20.10 39.26 24.18 42.09 17.65 44.19 24.54 46.78 36.32
-Table 8: Scores (%) of MLLMs on HSSBench (ZH-\@slowromancapii@).
-Figure 10 illustrates the performance of various models under four different prompt configurations.
-Figure 10: Comparison of Model Performances Across Four Prompt Settings.
-Table 9 and Figure 11 shows the performance of the model on datasets in different languages.
-Our dataset includes data organized in six languages.
-The table below presents the experimental results on a stratified sample of 900 instances.
-Model Arabic Chinese English French Russian Spanish
-InternVL3-8B 35.20 49.37 42.38 39.37 38.70 39.21
-Qwen2.5-VL-7B-Instruct 33.74 40.21 38.37 34.08 35.55 34.83
-Qwen2.5-VL-32B-Instruct 46.94 54.79 50.24 47.57 48.51 50.24
-Qwen2.5-VL-72B-Instruct 49.72 55.41 51.28 51.28 48.27 48.83
-QVQ-72B-Preview 33.98 36.89 39.80 37.86 40.77 38.83
-GPT-4.1-mini 41.33 44.89 41.67 41.78 41.44 46.89
-Average 40.82 46.93 43.96 41.99 42.87 43.47
-Table 9: Contrast Scores (%) of MLLMs on HSSBench for six UN languages and six models.
-Figure 11: Contrast Scores (%) of MLLMs on HSSBench for six UN languages and six models.
-C.4 Prompt for Model Inference
+C.5 Comparison with HSS-Related Benchmarks
+We acknowledge that some existing benchmarks, such as MME, include HSS-related test data, particularly within the Art domain.
+To better understand the relationship between these datasets and our proposed HSSBench, we conducted a detailed comparative analysis focusing on the overlapping Art category.
+The evaluation was performed under identical prompt settings, including both Direct and Chain-of-Thought (CoT) prompting.
+The combined results are presented in Table 13.
+Model Benchmark (% Accuracy) Prompting
+CMMMU MME MMMU HSSBench Direct CoT
+Qwen2.5-VL-3B-Instruct 47.73 77.32 57.76 29.01 47.73 48.86
+Qwen2.5-VL-7B-Instruct 43.18 71.21 61.21 37.88 43.18 46.59
+InternVL3-8B 65.91 85.28 68.10 42.14 65.91 55.68
+llava-onevision-qwen2-7b 53.41 77.21 49.14 36.20 53.41 52.27
+Table 13: Performance comparison on the Art category across HSS-related benchmarks under Direct and CoT prompting.
+As shown, the relative performance trends across these benchmarks are broadly consistent, indicating that the challenges within the Art domain are similar across datasets.
+Notably, HSSBench presents a more challenging evaluation, reflected by generally lower accuracy scores, which underscores its value in pushing forward research on Humanities and Social Sciences tasks.
+C.6 Comparison with STEM Benchmarks
+While our primary focus is on Humanities and Social Sciences (HSS), we also recognize the importance of situating HSSBench within the broader landscape of STEM benchmarks.
+To this end, we provide a comparative overview of model performance on several representative STEM benchmarks alongside HSSBench under the Chain-of-Thought (CoT) prompting setting.
+The results are summarized in Table 14.
+Model MMLU Pro GPQA Diamond SWE-bench Verified MATH- 500 AIME 2024 LiveCode Bench OpenCompass- Reasoning HSSBench
+Qwen2.5-VL-72B 71.2 50.2 51.87
+GPT-4.1 mini 65.00 23.60 49.60 46.0 48.03
+GPT-4o 79.80 66.90 35.80 54.8 46.88
+GPT-4.1 81.80 66.30 54.6 92.80 48.10 40.50 54.0 42.66
+InternVL3-8B 41.4 41.42
+GPT-4.1 nano 50.30 29.40 34.2 35.83
+Janus-Pro-7B 19.1 31.66
+Table 14: Model performance comparison on STEM benchmarks and HSSBench (CoT prompting).
+Notes: All STEM benchmark and OpenCompass results are sourced from official publications or model release notes.
+A dash (—) indicates that the model did not publicly report results for that benchmark.
+Minor discrepancies may exist due to testing variability but do not affect the overall trend.
+This comparison reveals that while models generally achieve higher accuracy on STEM benchmarks (e.g., MMLU Pro, MATH-500), their performance on HSSBench is comparatively lower.
+This gap highlights the unique challenges posed by HSS tasks and the necessity of dedicated benchmarks like HSSBench to drive progress in this domain.
+C.7 Retrieval-Augmented Generation Evaluation
+We conducted additional experiments integrating Retrieval-Augmented Generation (RAG) with several smaller MLLMs.
+The retrieval database was constructed from a general knowledge corpus comprising Wikipedia and publicly available documents related to HSS.
+Model performance was evaluated under two prompting strategies: direct prompting and CoT prompting.
+Table 15: Performance of models under direct prompting.
+Model Geography Art Culture Social Science History Economy Overall
+Without RAG
+Qwen2.5-VL-7B-Instruct 60.09 28.05 27.19 53.50 54.51 40.09 43.89
+Qwen2.5-VL-3B-Instruct 55.40 33.48 44.24 54.50 50.82 35.59 45.56
+InternVL3-8B 54.93 37.56 43.78 62.00 51.64 44.14 48.82
+llava-onevision-qwen2-7b 46.01 30.77 34.10 46.50 38.93 29.28 37.43
+With RAG
+Qwen2.5-VL-7B-Instruct 47.42 29.86 29.95 51.00 45.08 38.74 40.24
+Qwen2.5-VL-3B-Instruct 49.77 33.03 35.48 48.50 43.44 33.78 40.55
+InternVL3-8B 53.52 35.75 35.94 59.50 49.59 37.39 45.10
+llava-onevision-qwen2-7b 43.19 32.13 26.73 41.00 32.38 21.17 32.57
+Table 16: Performance of models under COT prompting.
+Model Geography Art Culture Social Science History Economy Overall
+Without RAG
+Qwen2.5-VL-7B-Instruct 55.40 25.79 27.65 51.00 47.54 39.64 41.08
+Qwen2.5-VL-3B-Instruct 51.17 30.32 35.48 49.00 46.31 32.43 40.70
+InternVL3-8B 53.52 34.39 40.55 57.00 51.64 38.74 45.86
+llava-onevision-qwen2-7b 43.66 31.22 28.57 43.00 33.61 29.73 34.78
+With RAG
+Qwen2.5-VL-7B-Instruct 47.89 27.15 27.19 46.50 42.62 36.49 37.89
+Qwen2.5-VL-3B-Instruct 41.78 31.22 29.95 44.00 40.57 36.04 37.21
+InternVL3-8B 57.75 30.77 34.10 55.50 50.41 39.19 44.50
+llava-onevision-qwen2-7b 42.25 29.86 28.11 42.50 31.56 23.42 32.73
+The results indicate that, although RAG occasionally yields modest improvements in specific domains or models, it does not consistently outperform direct prompting without retrieval augmentation.
+This suggests that augmenting MLLMs with a general retrieval corpus and straightforward prompting strategies may be insufficient to fully exploit the complex and nuanced knowledge required for HSS tasks.
+These findings highlight the challenges inherent in applying RAG to Humanities and Social Sciences benchmarks such as HSSBench.
+We hypothesize that more specialized, domain-specific retrieval corpora, combined with advanced retrieval and integration techniques, are necessary to unlock the full potential of RAG in this context.
+C.8 Prompt for Model Inference
 Table 17 details the configurations of the four prompts employed in our experiments, specifying the presence or absence of a CoT prompt and indicating whether the questions were open-ended or multiple-choice.
@@
 Table 17: Prompt for model inference.
-C.5 Evaluation for model’s output
+C.9 Evaluation for model’s output
 Table 18 presents the evaluation prompts employed to assess the accuracy of the model’s responses.
@@
 Appendix D More Cases
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4.
-Justification: We have detailed in the abstract and introduction that HSS-related issues still pose significant challenges for large models.
-Our contribution lies in proposing a large-scale, high-quality benchmark to evaluate the performance of large models on HSS tasks.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9.
-Justification: In Section 4, we discuss two limitations of our work: 1) Since there is no fixed set of classification tasks in the HSS field, our dataset does not cover all possible tasks within HSS; 2) Some data in certain HSS subfields have spatiotemporal relevance, so we can only ensure that the questions in our dataset are correct at the present time.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: Our work is focused on datasets and does not involve theoretical assumptions or proofs.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19.
-Justification: Our experimental setup prompts are described in detail in Appendix C.4 and  C.5.
-We also verify the consistency of evaluation across different models; that is, similar results can be achieved using other open-source models as with our closed-source model tests.
-Our work can be reproduced on our dataset from both the testing and evaluation perspectives.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: We have open-sourced our dataset and implementation code on both GitHub and Hugging Face.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: In Section 3.1, we provide a detailed description of our experimental design. Our experimental setup prompts are described in detail in Appendix C.4 and  C.5.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34.
-Justification: In the Appendix C.5, we discuss the consistency among human evaluation, GPT evaluation, and Qwen evaluation.
-The consistency is within an acceptable range, demonstrating that the choice of evaluation method does not have a significant statistical impact on the final results.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [N/A]
-  39. Justification: Our benchmark work is not related to computing resources; as long as it can run, that’s sufficient.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44.
-Justification: Our work complies with ethical standards, and all experts involved in data production are members of our team.
-Moreover, the images we use are either drawn by our experts or sourced from open-source data or websites.
-Our research does not have any harmful social impact.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49.
-Justification: Throughout this paper, we emphasize the positive social impact of our work.
-Our evaluation dataset can help the AI community assess the performance of large models on HSS tasks, while also highlighting the current insufficient attention given to HSS tasks.
-We believe our work can actively advance the development of MLLMs.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [Yes]
-  54.
-Justification: The construction of our dataset involves the use of multiple agents for web data crawling.
-However, we have implemented numerous restrictions during the crawling process, and image and text information can only be obtained from websites that we have selected and approved.
-We assure you that there are no risks associated with the use of our networked multi-agent system.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: We obtained data from https://www.unesco.org/zh/open-access/cc-nc-sa and comply with the CC BY-NC-SA license.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: Our dataset contains 13,152 image-text pairs, and each entry supports evaluation in the six official languages of the United Nations
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: Our dataset work does not involve crowdsourcing or research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: Our dataset work does not involve crowdsourcing or research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.